### v1

# Taming Rectified Flow for Inversion and Editing

Abstract
Rectified-flow-based diffusion transformers, such as FLUX and OpenSora, have demonstrated exceptional performance in the field of image and video generation. Despite their robust generative capabilities, these models often suffer from inaccurate inversion, which could further limit their effectiveness in downstream tasks such as image and video editing. To address this issue, we propose RF-Solver, a novel training-free sampler that enhances inversion precision by reducing errors in the process of solving rectified flow ODEs. Specifically, we derive the exact formulation of the rectified flow ODE and perform a high-order Taylor expansion to estimate its nonlinear components, significantly decreasing the approximation error at each timestep. Building upon RF-Solver, we further design RF-Edit, which comprises specialized sub-modules for image and video editing. By sharing self-attention layer features during the editing process, RF-Edit effectively preserves the structural information of the source image or video while achieving high-quality editing results. Our approach is compatible with any pre-trained rectified-flow-based models for image and video tasks, requiring no additional training or optimization. Extensive experiments on text-to-image generation, image & video inversion, and image & video editing demonstrate the robust performance and adaptability of our methods. Code is available at this URL.
{strip}[Uncaptioned image]
Figure 1: We propose RF-Solver to solve the rectified flow ODE with less error, thus enhancing both sampling quality and inversion-reconstruction accuracy for rectified-flow-based generative models. Furthermore, we propose RF-Edit to leverage the RF-Solver for image and video editing tasks. Our methods achieve impressive performance on various tasks, including text-to-image generation, image & video inversion, and image & video editing.
1 Introduction
Recently, rectified-flow-based generation methods [31] have demonstrated remarkable performance in generating high-quality images and videos. Compared with traditional approaches such as Stable Diffusion [19, 42], their impressive capabilities stem primarily from two aspects. First, rectified flow utilizes a straightforward Ordinary Differential Equation (ODE), which constructs a continuous straight-line motion system to produce the desired data distribution. This simplified training and inference paradigm enables the model to more effectively learn the underlying distribution of real image data. Second, most of the latest rectified flow methods employ the Diffusion Transformer (DiT) [38, 59] architecture as the backbone, which has shown superior performance compared to traditional U-Net [43] architectures. As a result, image generation models like FLUX [1] and video generation models like OpenSora [2], both based on rectified flow and DiT, have respectively become one of the state-of-the-art (SOTA) models in Text-to-Image (T2I) and Text-to-Video (T2V) generation.
Apart from fundamental T2I and T2V tasks, other downstream tasks such as reconstruction [46, 35, 16] and editing [17] have been attracting growing interest. These tasks typically depend on performing inversion on the source image/video. The inversion process yields the corresponding representation in the noise space, which is followed by denoising with various conditions provided by users. Given the robust generative capabilities of rectified-flow-based models, they are expected to exhibit superior performance on these downstream tasks. However, their performance remains unsatisfactory compared with traditional methods (based on DDPM and UNets), and research in this area still lags behind.
Delving into this problem, we identify that the primary challenge lies in the significant errors during the inversion process of the rectified flow, which fail to accurately reconstruct the original images or videos (Task 1 and Task 2 in Fig. 1). This limitation further constrains its performance on other downstream tasks such as editing. Although [45] attempted to address this challenge through dynamic optimal control, its application is restricted to simple image editing scenarios such as stylization and facial editing, which typically involve simple content and uniform backgrounds. In contrast, both the image contents and editing requirements (such as addition, replacement, and global editing) are considerably more complex in the real world. Moreover, video editing requires highly consistent temporal modeling, presenting even greater challenges for editing algorithms. To the best of our knowledge, there are still no existing methods that effectively tackle these issues.
Instead of focusing on designing a specific inversion method, we aim to address the above problem from a more general and fundamental perspective: the sampler. This is because the essence of the inversion and generation process for rectified flow is to employ a sampler that estimates the solution of rectified flow ODE. Consequently, the primary source of inaccurate inversion lies in the approximation error in the solution, which accumulates at each timestep. Intuitively, if the ODE is solved more accurately, the accuracy of inversion can be enhanced subsequently.
Based on this insight, we propose the RF-Solver. Specifically, we note that the ODE formulation for rectified flow can be solved directly using the variation of constants method, yielding an exact formulation of the solutions. For the nonlinear component of this solution (i.e., the integral of the neural network), we utilize Taylor expansion for estimation. By employing higher-order Taylor expansion, the ODE can be solved with reduced error, thereby improving the performance of rectified flow models. RF-Solver is a generic sampler that can be seamlessly integrated into any rectified flow model without the need for training or optimization. Experimental results demonstrate that RF-Solver not only significantly enhances the accuracy of inversion and reconstruction, but also improves performance on fundamental tasks such as T2I generation.
Building upon this, we propose RF-Edit to apply RF-Solver in editing tasks. Real-world image and video editing require the model to make precise modifications to a source image/video while maintaining its overall structure unchanged. This makes editing a more challenging task compared to reconstruction. In this scenario, it is inadequate to solely rely on the inverted noises as prior knowledge for editing, which could lead to edited results being excessively influenced by the target prompt, resulting in a completely different output compared with source image/video. Addressing this problem, we store the $V$ (value) feature in the self-attention layers at several timesteps during inversion. In the process of denoising, these features are used to replace the corresponding features. Practically, we design two specific sub-modules for RF-Edit, respectively leveraging FLUX [1] and OpenSora [2] as the backbone for image and video editing. With the effective design of RF-Edit, it demonstrates superior performance in both image and video domains, outperforming various SOTA methods.
Our core contributions are summarized as follows:
  • We propose RF-Solver, a training-free sampler that significantly reduces errors in the inversion and reconstruction processes of the rectified-flow model.
  • We present RF-Edit, which leverages RF-Solver for image and video editing, effectively preserving the structural integrity of the source image/video while achieving high-quality results.
  • Extensive experiments on images and videos demonstrate the efficacy of our methods, showcasing superior performance in both inversion and high-quality editing compared to various existing baselines.
2 Related Work
2.1 Inversion
Inversion maps the real visual data, i.e. image and video, to a representation in the noise space [14, 50, 35, 44, 34, 23], which is the reverse process of generation. This representation captures the essential features and structures of the original data while allowing for flexibility in manipulation for various editing applications. Numerous previous inversion approaches have been elaborately designed for diffusion models to achieve remarkable performance. The representative method, DDIM inversion [46, 47], adds predicted noise recursively at each forward step, outputting the final state as structured noise. To mitigate the discretization error in DDIM inversion, some efforts have been explored from the perspective of optimizing null prompt embeddings [35] and latent variables [44]. Negative prompt inversion [34] accelerates the inversion process at the cost of fidelity. Despite the success of inversion in diffusion models, the exploration of inversion in SOTA rectified flow models like FLUX and OpenSora is limited. RF-prior [57] uses the score distillation to invert the image while it requires a number of optimizing steps. More recently, [45] introduces an additional vector field that is conditioned on the source image to improve the inversion. However, its performance improvement mainly stems from the information leakage of the source image during the inversion and reconstruction process. The error from the original vector field of rectified flow still persists, which would limit the performance of such method on various downstream tasks.
2.2 Image and Video Editing
Training-free methods for image and video editing [21, 48] have gained increasing popularity due to their efficiency and effectiveness. These methods usually invert the source image/video into Gaussian noises and then denoises it conditioned on the target prompt. Existing image editing methods focus on prompt refinement [41, 52], attention-sharing mechanism [17, 37, 7, 49], mask guidance [5, 11, 20, 28], and noise initialization [6, 58]. Video editing introduces additional complexities due to the need for maintaining temporal consistency, making it a more challenging task. Existing video editing methods focus on attention injection [39, 53, 30], motion guidance [10, 15, 56, 51], latent manipulation [61, 55, 25, 9], and canonical representation [8, 27, 36, 26]. To date, the editing performance of rectified-flow-based diffusion transformers has remained largely under-explored. Although [45] employs FLUX for image editing, its performance is limited to simple tasks such as stylization and face editing, and it struggles to effectively maintain the structural information of source images. Moreover, currently there is no research exploring the video editing capabilities of rectified-flow-based models.
3 Method
In this section, we present our methods in detail. First, we introduce the proposed RF-Solver, which significantly enhances the precision of inversion and reconstruction. Subsequently, we introduce RF-Edit, an extension of RF-Solver designed to enable high-quality image and video editing.
3.1 Preliminaries
Rectified Flow [32] facilitates the transport between the Gaussian Noises distributions $\pi_{0}$ and real data distribution $\pi_{1}$ along a straight path. This is achieved by learning a forward-simulating system given by $\frac{d\boldsymbol{Z}_{t}}{dt}=v(\boldsymbol{Z}_{t},t),t\in[0,1]$ which maps $\boldsymbol{Z}_{0}\in\pi_{0}$ to $\boldsymbol{Z}_{1}\in\pi_{1}$. In practice, the velocity field $v$ is parameterized by a neural network $\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t},t)$. During training, given $\boldsymbol{X}_{0}\in\pi_{0}$ and $\boldsymbol{X}_{1}\in\pi_{1}$, the forward process of rectified flow is implemented through a simple linear combination:
$\boldsymbol{X}_{t}=t\boldsymbol{X}_{1}+(1-t)\boldsymbol{X}_{0},t\in[0,1].$ (1)
The rectified flow ODE can be derived by differentiating Eq. 1 with respect to $t$, as follows:
$\frac{d\boldsymbol{X}_{t}}{dt}=\boldsymbol{X}_{1}-\boldsymbol{X}_{0},t\in[0,1].$ (2)
Here, $\boldsymbol{X}_{1}-\boldsymbol{X}_{0}$ serves as the ground truth, and the network is optimized through mean square error:
[…]
During the sampling process, the ODE is discretized and solved using Euler method. Specifically, the rectified flow model starts with Gaussian noises $\boldsymbol{Z}_{t_{N}}\in\mathcal{N}(0,\boldsymbol{I})$. Given a series of discrete $N$ timesteps $t=\{t_{N},...,t_{0}\}$, the model iteratively predicts $\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}}),i\in\{N,\cdots,1\}$ and then takes a step forward until generating the images $\boldsymbol{Z}_{t_{0}}$, with the following recurrence relation:
$\boldsymbol{Z}_{t_{i-1}}=\boldsymbol{Z}_{t_{i}}+(t_{i-1}-t_{i})\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i}).$ (4)
The definition and training process of Rectified Flow is designed to establish a nearly linear transition trajectory between two distributions, enabling efficient generation with significantly fewer steps compared to DDPMs [19].
3.2 RF-Solver
Figure 2: Analysis of the Inversion-Reconstruction Process. Inversion begins with the source image latent $\widetilde{\boldsymbol{{Z}}}_{t_{0}}$ as the input and progressively add noise, obtaining $\widetilde{\boldsymbol{{Z}}}_{t_{N}}\in\mathcal{N}(0,\boldsymbol{I})$. $\widetilde{\boldsymbol{Z}}_{t_{N}}$ is then denoised for $N$ timesteps to obtain the reconstruction $\boldsymbol{Z}_{t_{0}}$. During this process, we store the latent $\boldsymbol{\widetilde{Z}}_{t_{i}}$ and $\boldsymbol{Z}_{t_{i}}$ at each time step and calculate the mean squared error between them.
Figure 3: RF-Edit Pipelines. The RF-Edit framework comprises two sub-modules, respectively applied to FLUX [1] for image editing and OpenSora [2] for video editing. For image editing, we share the feature of $\widetilde{\mathcal{V}}$ within the single block self-attention of the FLUX [1] backbone. For video editing, we share the feature of $\widetilde{\mathcal{V}}$ in the spatial attention of the OpenSora [2] backbone.
The vanilla rectified flow (RF) sampler demonstrates strong performance in image and video generation. However, when applied to inversion and reconstruction tasks, we observe significant error accumulation at each timestep. This results in reconstructions that diverge notably from the original image (see Fig. 2), further limiting the performance of RF-based models in various downstream tasks, such as image and video editing. Delving into this problem, we notice that the inversion and reconstruction processes in rectified flow rely on estimating an approximate solution of the rectified flow ODE at each timestep (see Eq. 4). Obtaining more precise solutions for the ODE would effectively mitigate these errors, leading to improved reconstruction quality. Based on this analysis, we start by carefully examining the differential form of the Rectified flow: $\frac{d\boldsymbol{Z}_{t}}{dt}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t},t)$. This ODE is discretized in the sampling process. Given the initial value $\boldsymbol{Z}_{t_{i}}$, the ODE can be exactly formulated using the variant of constant method:
[…]
In the above formula, $\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{\tau},\tau)$ is the non-linear component parameterized by the complex neural network, which is difficult to approximate directly. As an alternative, we employ the Taylor expansion at ${t_{i}}$ to approximate this term:
$\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{\tau},\tau)=\sum\limits_{k=0}^{n-1}\frac{(\tau-{t_{i}})^{k}}{k!}\boldsymbol{v}_{\theta}^{(k)}(\boldsymbol{Z}_{t_{i}},{t_{i}})+\mathcal{O}\big{(}(\tau-{t_{i}})^{n}\big{)},$ (6)
where $\boldsymbol{v}^{(k)}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i})=\frac{\text{d}^{k}\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i})}{\text{d}t^{k}}$, denoting the $k$-order derivative of $\boldsymbol{v}_{\theta}$ and $\mathcal{O}$ denotes higher-order infinitesimals. Substituting Eq. 6 into the integral term yields:
[…]
$\quad+\mathcal{O}\big{(}(\tau-{t_{i}})^{n}\big{)}.$ (7)
Through the above process, the network prediction term and its higher-order derivatives are separated from the integral. Then we notice that the remaining portion in the integral can be computed analytically:
[…]
Substituting Eq. 8 and Sec. 3.2 into Eq. 5, we derive the $n$-th order solution of Rectified flow ODE:
$\boldsymbol{Z}_{t_{i-1}}=\boldsymbol{Z}_{t_{i}}$ $+\sum\limits_{k=0}^{n-1}\frac{({t_{i-1}}-{t_{i}})^{k+1}}{(k+1)!}{\boldsymbol{v}}^{(k)}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$
$+\mathcal{O}\big{(}h_{i}^{n+1}\big{)},$ (9)
where $h_{i}:=t_{i-1}-t_{i}$. Sec. 3.2 indicates that to estimate $\boldsymbol{Z}_{t_{i-1}}$, we need to obtain the $k$-th order derivatives $\{\boldsymbol{v}^{(k)}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})\}$ for $k\in\{0,\cdots,n-1\}$.
Algorithm 1 Sampling process of RF-Solver
Input:
  $\boldsymbol{v}_{\theta}$triangleright Velocity function
  $t=[t_{N},\ldots,t_{0}]$triangleright Time steps
  $\boldsymbol{Z}_{t_{N}}\sim\mathcal{N}(0,I)$triangleright Initial Gaussian Noise
For $i=N$ to $1$ do
  $\Delta t_{i}\leftarrow\frac{1}{2}(t_{i-1}-t_{i})$
  $\boldsymbol{\hat{v}}_{t_{i}}\leftarrow\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i})$
  $\boldsymbol{Z}_{t_{i}+\Delta t_{i}}\leftarrow\boldsymbol{Z}_{t_{i}}+\Delta t_{i}\boldsymbol{\hat{v}}_{t_{i}}$
  $\boldsymbol{\hat{v}}_{t_{i}+\Delta t_{i}}\leftarrow\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}+\Delta t_{i}},t_{i}+\Delta t_{i})$
  $\boldsymbol{v}_{t_{i}}^{(1)}\leftarrow({\boldsymbol{\hat{v}}_{t_{i}}-\boldsymbol{\hat{v}}_{t_{i}+\Delta t_{i}}})/{\Delta t_{i}}$triangleright Calculating the Derivatives
  $\boldsymbol{Z}_{t_{i-1}}\leftarrow\boldsymbol{Z}_{t_{i}}+(t_{i-1}-t_{i})\boldsymbol{\hat{v}}_{t_{i}}+\frac{1}{2}(t_{i-1}-t_{i})^{2}\boldsymbol{v}_{t_{i}}^{(1)}$
Output: $\boldsymbol{Z}_{0}$
When $n=1$, the formula reduces to the standard rectified flow (Eq. 4). In our experiments, we find that setting $n=2$ effectively mitigates the errors, yielding the following formula:
$\boldsymbol{Z}_{t_{i-1}}=\boldsymbol{Z}_{t_{i}}$ $+({t_{i-1}}-{t_{i}})\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i})$
$+\frac{1}{2}(t_{i-1}-t_{i})^{2}\boldsymbol{v}_{\theta}^{(1)}(\boldsymbol{Z}_{t_{i}},t_{i}).$ (10)
Note that $\boldsymbol{v}_{\theta}^{(1)}$ is the first-order derivative of the network prediction term $\boldsymbol{v}_{\theta}$, which cannot be analytically derived due to the complex architecture of the neural network. To estimate this term, we first obtain the network prediction $\hat{\boldsymbol{v}}_{t_{i}}$ at the timestep $t_{i}$, i.e., $\hat{\boldsymbol{v}}_{t_{i}}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$. Then we step forward a small timestep $\Delta t=\frac{1}{2}(t_{i-1}-t_{i})$, and update the latents to obtain $\boldsymbol{Z}_{{t_{i}}+\Delta t}=\boldsymbol{Z}_{t_{i}}+\Delta t\cdot\hat{\boldsymbol{v}}_{t_{i}}$. Next, we calculate an additional prediction of the network at the timestep $t_{i}+\Delta t$, i.e., $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}+\Delta t},{t_{i}+\Delta t})$. With $\hat{\boldsymbol{v}}_{t_{i}}$ and $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}$, the first-order derivative of ${\boldsymbol{v}}_{\theta}$ at the timestep $t_{i}$ can be estimated as:
$\boldsymbol{v}_{\theta}^{(1)}(\boldsymbol{Z}_{t_{i}},t_{i})=\frac{\hat{\boldsymbol{v}}_{t_{i}+\Delta t}-\hat{\boldsymbol{v}}_{t_{i}}}{\Delta t}.$ (11)
Substituting Eq. 11 into Sec. 3.2 results in the practical implementation of the RF-Solver algorithm. The complete sampling process for RF-Solver is presented in Algorithm 1.
Besides sampling, inversion seeks to map data back into noise, which reverses the sampling process. Following previous methods for DDIM inversion [46, 13], the ODE process can be directly reversed in the limit of small steps. Based on this assumption, the inversion process of RF-Solver (Sec. 3.2) can be directly transformed as:
$\widetilde{\boldsymbol{Z}}_{t_{i+1}}=\widetilde{\boldsymbol{Z}}_{t_{i}}$ $+({t_{i+1}}-{t_{i}})\boldsymbol{v}_{\theta}(\widetilde{\boldsymbol{Z}}_{t_{i}},t_{i})$
$+\frac{1}{2}(t_{i+1}-t_{i})^{2}\boldsymbol{v}_{\theta}^{(1)}(\widetilde{\boldsymbol{Z}}_{t_{i}},t_{i}),$ (12)
where $\widetilde{\boldsymbol{Z}}_{t_{i}}$ and $\widetilde{\boldsymbol{Z}}_{t_{i+1}}$ denotes the latents during inversion. Through this high order expansion, the error of the ODE solution in each timestep is reduced from $\mathcal{O}\big{(}(h_{i})^{2}\big{)}$ to $\mathcal{O}\big{(}(h_{i})^{3}\big{)}$, leading to improved performance, particularly in inversion and reconstruction (see Fig. 2). Beyond inversion and reconstruction, RF-Solver can also be applied to any RF-based model (such as FLUX [1] and OpenSora [2]) for other tasks such as sampling and editing, enhancing performance without requiring additional training.
3.3 RF-Edit
Incorporating higher-order terms enables RF-Solver to significantly reduce errors in the ODE-solving process, improving both sampling quality and inversion accuracy. Furthermore, we extend the application of RF-Solver to real-world image and video editing tasks, which are more challenging than reconstruction. In these scenarios, maintaining the content and structure of the original image is crucial. For instance, when adding new objects to a source image, other unrelated regions should remain unaffected by the editing process. However, directly applying RF-Solver during the inversion and denoising stages may cause the model to be overly influenced by the target prompt, leading to unintended modifications in other parts of the image or video (e.g., altering the unrelated which is not mentioned in the editing prompt). This issue is common across various existing editing methods [45, 17, 49].
To address this problem, we propose RF-Edit, which incorporates features from the inversion process into the denoising procedure. Specifically, during the last $n$ steps of inversion, we extract and store the Value feature $\{\mathcal{{\widetilde{V}}}_{t_{k}}^{m}\}$ and $\{\mathcal{{\widetilde{V}}}_{t_{k}+\Delta t_{k}}^{m}\}$ from the self-attention layers in the last $M$ transformer blocks at each timestep $k$. Here, $k\in\{n,n+1,\cdots,N\}$ and $m\in\{1,\cdots,M\}$. This process can be formulated as follows:
$\{\mathcal{{\widetilde{V}}}_{t_{k}}^{m}\}$ $=\text{Extract}\big{(}\boldsymbol{v}_{\theta}(\boldsymbol{\widetilde{Z}}_{t_{k}},t_{k})\big{)}$ (13)
$\{\mathcal{{\widetilde{V}}}^{m}_{t_{k}+\Delta t_{k}}\}$ $=\text{Extract}\big{(}\boldsymbol{v}_{\theta}(\boldsymbol{\widetilde{Z}}_{t_{k}+\Delta t_{k}},t_{k}+\Delta t_{k})\big{)},$ (14)
where $\boldsymbol{\widetilde{Z}}_{t_{k}}$ and $\boldsymbol{\widetilde{Z}}_{t_{k}+\Delta t_{k}}$ denote the latents during inversion for RF-Solver.
During the first $n$ timesteps of denoising, considering the $m$th transformer block at the timestep $k$, the original self-attention in the network $\boldsymbol{v}_{\theta}$ can be formulated as:
$\boldsymbol{F}^{m}_{t_{k}}=\text{Attention}(\mathcal{{Q}}_{t_{k}}^{m},\mathcal{{K}}_{t_{k}}^{m},\mathcal{{V}}_{t_{k}}^{m}),$ (15)
where $\boldsymbol{F}^{m}_{t_{k}}$ denotes the output feature of the self-attention module and $\mathcal{{Q}}_{t_{k}}^{m},\mathcal{{K}}_{t_{k}}^{m},\mathcal{{V}}_{t_{k}}^{m}$ represent query, key and value for attention during the denoising process, respectively.
Figure 4: Qualitative Results of Image and Video Reconstruction. Our method (The second row) illustrates better performance compared with the rectified flow baselines (The third row) on both image and video reconstruction.
In RF-Edit, the above self-attention mechanism is modified to cross-attention where $\mathcal{{V}}_{t_{k}}^{m}$ is replaced by $\mathcal{{\widetilde{V}}}_{t_{k}}^{m}$,
${{\boldsymbol{F}}^{m}_{t_{k}}}^{\prime}=\text{Attention}(\mathcal{{Q}}_{t_{k}}^{m},\mathcal{{K}}_{t_{k}}^{m},\mathcal{{\widetilde{V}}}_{t_{k}}^{m}).$ (16)
The modified output feature ${{\boldsymbol{F}}^{m}_{t_{k}}}^{\prime}$ is then passed to the subsequent modules for further processing.
Similarly, this feature-sharing process is also adopted in the derivative calculation process of RF-Solver:
${{\boldsymbol{F}}^{m\prime}_{t_{k+\Delta t_{k}}}}=\text{Attention}(\mathcal{{Q}}_{t_{k+\Delta t_{k}}}^{m},\mathcal{{K}}_{{k+\Delta t_{k}}}^{m},\mathcal{{\widetilde{V}}}_{{k+\Delta t_{k}}}^{m}).$ (17)
The proposed RF-Edit framework enables high-quality editing while preserving structural information. Building on this concept, we design two sub-modules for RF-Edit, specifically tailored for image editing and video editing. For image editing, we use FLUX [1] as the backbone, which comprises several double blocks and single blocks. Double blocks independently modulate text and image features, while single blocks concatenate these features for unified modulation. In this architecture, RF-Edit shares features within the single blocks, as they capture information from both the source image and the source prompt, enhancing the ability of the model to preserve the structural information of the source image. For video editing, we employ OpenSora [2] as the backbone. The DiT blocks in OpenSora include spatial attention, temporal attention, and text cross-attention. Within this architecture, the structural information of the source video is captured in the spatial attention module, where we implement feature sharing.
4 Experiment
4.1 Setup
Baselines. We select the vanilla Rectified Flow sampler as the primary baseline for all tasks. Additionally, for image editing, we compare our method with P2P [17], DiffEdit [12], SDEdit [33], PnP [49], Pix2pix [37] and RF-Inversion [45]. For video editing tasks, we compare our methods with FateZero [39], FLATTEN [10], COVE [51], RAVE [25], Tokenflow [15]. Detailed experimental settings of these methods are provided in the Appendix.
Figure 5: Qualitative Results of Text-to-Image Generation. With the RF-Solver, the model can generate images with higher quality (The first row) than baselines (The Second row).
Metric FID ($\downarrow$) Clip Score ($\uparrow$)
RF 26.55 31.01
Ours 25.45 31.09
Table 1: Quantitive results on Text-to-Image Generation. RF-Solver outperforms the vanilla RF sampler.
Figure 6: Qualititive Comparison of Image Editing. With RF-Solver and feature-sharing mechanism in RF-Edit, our method can successfully handle various kinds of image editing cases, outperforming the previous SOTA methods. Zoom-in for the best views.
Implementation Details. In the experiment, we adopt the guidance-distilled variant of FLUX [1] for image tasks and OpenSora [2] for video tasks. For sampling, we set the denoising step to 10 for our method. For reconstruction and editing, the inversion and denoising steps are set to 30. The derivative computation in RF-Solver requires an additional forward pass, resulting in the network needing to forward twice at each timestep. As a result, when comparing our method with the Rectified Flow baselines, we set the number of timestep for the vanilla Rectified Flow to be twice that of our method in order to maintain the same number of function evaluations (NFE) to ensure a fair comparison. Both the baseline and our method are conducted on a single A100 GPU with 40GB memory.
Evaluation Metrics For text-to-image sampling, we randomly select 10k images from the MSCOCO validation set [29] and report the FID [18] and Clip Score [40]. For the inversion and reconstruction task, we report the Mean Square Error (MSE), LPIPS [60], SSIM [54], and PSNR [24]. For image editing tasks, we report the Clip Score [40], which reflects whether the edited images align with the target prompt, and LPIPS [60], which reflects whether the edited images preserve the content of the source images. For video editing tasks, we adopt the metric proposed by [22], including Subject Consistency (SC), Motion Smoothness (MS), Aesthetic Quality (AQ), and Imaging Quality (IQ). SC and MS assess the temporal consistency of the edited video, while AQ and IQ assess the visual quality.
4.2 Text-to-image Sampling
We compare the performance of our method with the vanilla rectified flow on the text-to-image generation task. Both the quantitative (Tab. 1) and qualitative results (Fig. 5) demonstrate the superior performance of RF-Solver in fundamental T2I generation tasks, producing higher-quality images that align more closely with human cognition.
4.3 Inversion and Reconstruction
We conduct experiments on inversion and reconstruction for both image and video modalities, comparing our methods with the vanilla rectified flow sampler. For image inversion and reconstruction, we use images from the MSCOCO dataset, and for video inversion and reconstruction, we select videos from social media platforms such as TikTok and other publicly available sources [3, 4]. We use GPT-4.0 to annotate the content of the images and videos in detail, and then manually polish the GPT-generated content. These annotations are used as the source and target prompts for inversion and reconstruction.
Quantitive Comparison. The quantitative comparisons (Tab. 2) are conducted on both images and videos to illustrate the similarity between the source and reconstruction results. Our method demonstrates superior performance across all four metrics compared with vanilla rectified flow.
Mehtod MSE ($\downarrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$)
image RF 0.0268 0.6253 0.7626 28.28
Ours 0.0094 0.4242 0.9271 29.83
video RF 0.0206 0.4159 0.8134 18.12
Ours 0.0139 0.3299 0.8805 18.32
Table 2: Quantitative Results on Inversion and Reconstruction. Our methods significantly improve the reconstruction accuracy of both images and videos.
Qualitative Comparison. RF-Solver effectively reduces the error in the solution of RF ODE, thereby increasing the accuracy of the reconstruction. As illustrated in Fig. 4(a), the image reconstruction results using vanilla rectified flow exhibit noticeable drift from the source image, with significant alterations to the appearance of subjects such as cake, church, train, and jeep. For video reconstruction, as shown in Fig. 4(b), the baseline reconstruction results suffer from distortion. In contrast, RF-Solver significantly alleviate these issues, achieving more satisfactory results.
Figure 7: Qualititive Comparison of Video Editing. The first video has 200 frames with a resolution of $512\times 512$, and the second video has 60 frames with a resolution of $1024\times 768$ (we compress the frames for a neat layout). Our method outperforms various previous SOTA video editing methods, especially excelling in dealing with complex prompts.
4.4 Editing
We conduct experiments to evaluate the image and video editing performance of our methods. Image editing usually involves replacing the subject in the image with another one, adding new items, and global editing. For the first two types of editing, the background of the source image is expected to remain unchanged after editing. For global editing such as style transfer, the overall structure of the source image is expected to remain unchanged. Compared to image editing, video editing is much more challenging due to the complexity of modeling temporal motions in videos. Recent mainstream video editing methods usually focus on replacing the subjects and performing global editing.
Quantitative Comparison. For image editing, we perform quantitative comparison between our methods and baselines (Tab. 3), reporting the Clip score and LPIPS. Our method outperforms all other methods in Clip score, indicating that the edited images align well with the user-provided prompts. For LPIPS, it is noted that PnP [49] has a much lower value than all other methods. Based on the qualitative results (Fig. 6), it can be seen that PnP is only suitable for editing cases that do not significantly modify the structure or shape of the source image (such as changing red roses into yellow sunflowers). For shape editing, such as modifying a car into a motorbike, PnP fails to edit this case, resulting in an image very similar to the source. Consequently, although PnP has the lowest LPIPS score, its clip score is the lowest.
P2P DiffEdit SDEdit PnP Pix2Pix RF-Inv Ours
LPIPS ($\downarrow$) 0.419 0.157 0.394 0.080 0.155 0.318 0.149
Clip Score ($\uparrow$) 30.70 32.68 31.61 30.58 32.33 33.02 33.66
Table 3: Quantitative Results of Image Editing. RF-Edit can effectively edit the images according to the prompts while keeping the unrelated regions unchanged.
For video editing, we compare our methods with baseline methods using the VBench [22] metrics (Tab. 4). The results illustrate that our methods successfully maintain temporal consistency in long videos, achieving the highest Subject Consistency (SC) and Motion Smoothness (MS) scores. Additionally, our method demonstrates superior visual quality, outperforming the baselines at Aesthetic Quality (AQ) and Image Quality (IQ).
FateZero Flatten COVE RAVE Tokenflow Ours
SC ($\uparrow$) 0.9382 0.9420 0.9433 0.9292 0.9439 0.9501
MS ($\uparrow$) 0.9611 0.9528 0.9697 0.9519 0.9632 0.9712
AQ ($\uparrow$) 0.6092 0.6329 0.6717 0.6586 0.6742 0.6796
IQ ($\uparrow$) 0.6898 0.7024 0.7163 0.6917 0.7128 0.7207
Table 4: Quantitative Results of Video Editing. RF-Edit outperforms a number of previous SOTA video editing methods.
Qualitative Comparison. For image editing, we compare the performance of our methods with several baselines across different types of editing tasks (Fig. 6). The baseline methods often suffer from background changes or fail to perform the desired edits. In contrast, our methods demonstrate satisfying performance, effectively achieves a balanced trade-off between the fidelity to the target prompt and preservation of the source image. To be noticed, although RF-inversion [45] also uses the rectified flow model for image editing (third row in Fig. 6), the structure of the source image which is unrelated to editing prompt (such as background and human appearance) is modified obviously.
For video editing, we primarily evaluate the performance of our methods on long videos (200 frames) and high-resolution videos ($1280\times 768$). Furthermore, we assess the performance on complicated videos and prompts where there are multiple objects in the video, and the user has different editing requirements for each object. The qualitative results are shown in Fig. 7. Our method successfully handles complicated editing cases (e.g., modifying the leftmost lion among three lions into a white polar bear and changing the other two small lions into orange tiger cubs), whereas all other baseline methods fail in this scenario. Our method also demonstrates strong performance in global editing tasks, such as transforming scenes into autumn.
4.5 Ablation Study
We conduct ablation studies to illustrate the effectiveness of RF-Solver and RF-Edit. Without loss of generality, these ablation studies are performed on the image tasks using FLUX [1] as the base model.
Taylor Expansion Order of RF-Solver. We investigated the impact of the Taylor expansion order in RF-Solver (Tab. 5) under the same NFE across different orders. The second-order expansion demonstrated a significant improvement across various tasks compared to the first-order expansion (i.e., the vanilla rectified flow). However, higher-order expansions did not yield further enhancements. We speculate that this is primarily due to higher-order Taylor expansions requiring more inference steps per timestep. With a fixed NFE, this results in a reduced overall number of timesteps compared to lower-order expansions, leading to suboptimal performance. Moreover, computing the higher-order derivatives of $v_{\theta}{t_{i}}$ substantially increases the complexity of the algorithm, posing challenges for practical applications. Consequently, we predominantly employed second-order expansion in our experiments.
Metric RF RF-Solver-2 RF-Solver-3
Sampling FID ($\downarrow$) 26.55 25.45 25.37
Clip Score ($\uparrow$) 31.01 31.09 31.09
Inversion MSE ($\downarrow$) 0.0268 0.0094 0.0131
LPIPS ($\downarrow$) 0.6253 0.4242 0.4817
Editing LPIPS ($\downarrow$) 0.1524 0.1494 0.1503
Clip Score ($\uparrow$) 32.97 33.66 33.18
Table 5: Ablation Study on the Taylor Expansion Order. We choose the 2-order expansion (i.e. RF-Solver-2) for the downstream tasks for its effectiveness and simplicity.
Figure 8: Ablation Study of Feature-Sharing Step in RF-Edit. A too-small feature-sharing step leads to inconsistency between source and target images. On the other hand, a too-large feature-sharing step can cause the failure of editing.
Feature Sharing Steps of RF-Edit. RF-Edit leverages feature sharing to maintain the structural consistency between original images and edited images. However, an excessive number of feature-sharing steps may result in the edited output being overly similar to the source image, ultimately undermining the intended editing objectives (Fig. 8). To investigate the impact of feature-sharing steps on editing results, we incrementally increase the number of feature-sharing steps applied to the same image. Due to the varying levels of difficulty that different images presented to the model, the optimal number of sharing steps may differ across cases. Experimental results reveal that setting the sharing step to 5 effectively meets the editing requirements for most images. Additionally, we can customize the sharing step for each image to identify the most satisfying outcome.
5 Conclusion
In this paper, we propose RF-Solver, a versatile sampler for the rectified flow model that solves the rectified flow ODE with reduced error, thus enhancing the image and video generation quality across various tasks such as sampling and reconstruction. Based on RF-Solver, we further propose RF-Edit, which achieves high-quality editing performance while effectively preserving the structural information in source images or videos. Extensive experiments demonstrate the versatility and effectiveness of our methods.

### v2

@@
 Abstract
-Rectified-flow-based diffusion transformers, such as FLUX and OpenSora, have demonstrated exceptional performance in the field of image and video generation.
-Despite their robust generative capabilities, these models often suffer from inaccurate inversion, which could further limit their effectiveness in downstream tasks such as image and video editing.
-To address this issue, we propose RF-Solver, a novel training-free sampler that enhances inversion precision by reducing errors in the process of solving rectified flow ODEs.
-Specifically, we derive the exact formulation of the rectified flow ODE and perform a high-order Taylor expansion to estimate its nonlinear components, significantly decreasing the approximation error at each timestep.
-Building upon RF-Solver, we further design RF-Edit, which comprises specialized sub-modules for image and video editing.
-By sharing self-attention layer features during the editing process, RF-Edit effectively preserves the structural information of the source image or video while achieving high-quality editing results.
+Rectified-flow-based diffusion transformers like FLUX and OpenSora have demonstrated outstanding performance in the field of image and video generation.
+Despite their robust generative capabilities, these models often struggle with inversion inaccuracies, which could further limit their effectiveness in downstream tasks such as image and video editing.
+To address this issue, we propose RF-Solver, a novel training-free sampler that effectively enhances inversion precision by mitigating the errors in the ODE-solving process of rectified flow.
+Specifically, we derive the exact formulation of the rectified flow ODE and apply the high-order Taylor expansion to estimate its nonlinear components, significantly enhancing the precision of ODE solutions at each timestep.
+Building upon RF-Solver, we further propose RF-Edit, a general feature-sharing-based framework for image and video editing.
+By incorporating self-attention features from the inversion process into the editing process, RF-Edit effectively preserves the structural information of the source image or video while achieving high-quality editing results.
 Our approach is compatible with any pre-trained rectified-flow-based models for image and video tasks, requiring no additional training or optimization.
-Extensive experiments on text-to-image generation, image & video inversion, and image & video editing demonstrate the robust performance and adaptability of our methods.
-Code is available at this URL.
+Extensive experiments across generation, inversion, and editing tasks in both image and video modalities demonstrate the superiority and versatility of our method.
+The source code is available at this URL.
 {strip}[Uncaptioned image]
-Figure 1: We propose RF-Solver to solve the rectified flow ODE with less error, thus enhancing both sampling quality and inversion-reconstruction accuracy for rectified-flow-based generative models.
-Furthermore, we propose RF-Edit to leverage the RF-Solver for image and video editing tasks.
-Our methods achieve impressive performance on various tasks, including text-to-image generation, image & video inversion, and image & video editing.
+Figure 1: RF-Solver for downstream tasks in image and video.
+We propose RF-Solver to solve the rectified flow ODE with reduced error, thereby enhancing both sampling quality and inversion-reconstruction accuracy for rectified-flow-based generative models [2, 3].
+Furthermore, we propose RF-Edit, which utilizes the RF-Solver for image and video editing tasks.
+Our methods demonstrate impressive performance across generation, inversion, and editing tasks in both image and video modalities.
 1 Introduction
-Recently, rectified-flow-based generation methods [31] have demonstrated remarkable performance in generating high-quality images and videos.
-Compared with traditional approaches such as Stable Diffusion [19, 42], their impressive capabilities stem primarily from two aspects.
-First, rectified flow utilizes a straightforward Ordinary Differential Equation (ODE), which constructs a continuous straight-line motion system to produce the desired data distribution.
-This simplified training and inference paradigm enables the model to more effectively learn the underlying distribution of real image data.
-Second, most of the latest rectified flow methods employ the Diffusion Transformer (DiT) [38, 59] architecture as the backbone, which has shown superior performance compared to traditional U-Net [43] architectures.
-As a result, image generation models like FLUX [1] and video generation models like OpenSora [2], both based on rectified flow and DiT, have respectively become one of the state-of-the-art (SOTA) models in Text-to-Image (T2I) and Text-to-Video (T2V) generation.
-Apart from fundamental T2I and T2V tasks, other downstream tasks such as reconstruction [46, 35, 16] and editing [17] have been attracting growing interest.
-These tasks typically depend on performing inversion on the source image/video.
-The inversion process yields the corresponding representation in the noise space, which is followed by denoising with various conditions provided by users.
-Given the robust generative capabilities of rectified-flow-based models, they are expected to exhibit superior performance on these downstream tasks.
-However, their performance remains unsatisfactory compared with traditional methods (based on DDPM and UNets), and research in this area still lags behind.
-Delving into this problem, we identify that the primary challenge lies in the significant errors during the inversion process of the rectified flow, which fail to accurately reconstruct the original images or videos (Task 1 and Task 2 in Fig.
-1).
-This limitation further constrains its performance on other downstream tasks such as editing.
-Although [45] attempted to address this challenge through dynamic optimal control, its application is restricted to simple image editing scenarios such as stylization and facial editing, which typically involve simple content and uniform backgrounds.
-In contrast, both the image contents and editing requirements (such as addition, replacement, and global editing) are considerably more complex in the real world.
-Moreover, video editing requires highly consistent temporal modeling, presenting even greater challenges for editing algorithms.
-To the best of our knowledge, there are still no existing methods that effectively tackle these issues.
-Instead of focusing on designing a specific inversion method, we aim to address the above problem from a more general and fundamental perspective: the sampler.
-This is because the essence of the inversion and generation process for rectified flow is to employ a sampler that estimates the solution of rectified flow ODE.
-Consequently, the primary source of inaccurate inversion lies in the approximation error in the solution, which accumulates at each timestep.
-Intuitively, if the ODE is solved more accurately, the accuracy of inversion can be enhanced subsequently.
-Based on this insight, we propose the RF-Solver.
-Specifically, we note that the ODE formulation for rectified flow can be solved directly using the variation of constants method, yielding an exact formulation of the solutions.
+Recent advancements of generation methods based on Rectified Flow (RF) [42, 35, 67] have demonstrated exceptional performance in synthesizing high-quality images and videos.
+Different from traditional approaches represented by Stable Diffusion [22, 56], these methods leverage the Diffusion Transformer [51, 78, 73, 63] architecture and implement a straight-line motion system to produce the desired data distribution.
+With these effective designs, FLUX [2] and OpenSora [3] have respectively emerged as one of the state-of-the-art (SOTA) methods in the field of Text-to-Image (T2I) and Text-to-Video (T2V) generation.
+Despite the remarkable success in the fundamental T2I and T2V generation tasks, few studies have explored the performance of RF-based models on various downstream tasks such as inversion-reconstruction [60, 47, 17, 71] and editing [20, 45].
+When directly applying the vanilla RF for inversion, we observe that it fails to faithfully reconstruct the image or video from the source.
+Examples are shown in Fig.
+1 Task 1 and Task 2 (the third row).
+For image inversion, the positions of objects (e.g.
+the cake) and the appearance of individuals (e.g.
+the child) in the reconstructed image significantly diverge from the source image.
+The performance of video inversion is even worse, with noticeable distortions present in the reconstructed video.
+The inaccuracies of inversion and reconstruction would severely constrain the performance of RF models on other inversion-based downstream tasks such as image editing [20, 64, 48, 13, 29] and video editing [41, 33, 15, 59].
+In this work, we investigate the aforementioned problem by delving into the inversion and reconstruction process of the RF.
+Specifically, we track the latent at each intermediate timestep during inversion and reconstruction, calculating the Mean Square Error (MSE) between them at corresponding timesteps.
+We observe that significant errors are introduced at each timestep throughout the whole reconstruction process, and their accumulation ultimately results in a considerably deviated output (the red curve in Fig.
+2).
+Based on the definition and inference process of RF [40, 42], we identify that these errors stem from the Ordinary Differential Equation (ODE) solving process [79, 66, 23, 52].
+Specifically, the essence of the inversion and generation process for RF is to derive the solution of RF ODE [42].
+Since this ODE includes terms involving complex neural networks, the solution can only be coarsely approximated by a sampler.
+However, the experiment in Fig.
+2 indicates that the sampler adopted in existing models [2, 3] lacks sufficient precision for the inversion task, causing notable errors to accumulate at each timestep, finally leading to unsatisfactory reconstruction results.
+Figure 2: Analysis of the inversion-reconstruction process.
+Inversion takes the source image latent $\widetilde{\boldsymbol{{Z}}}_{t_{0}}$ as the input and progressively add noise for $N$ timesteps, obtaining $\widetilde{\boldsymbol{{Z}}}_{t_{N}}\in\mathcal{N}(0,\boldsymbol{I})$.
+$\widetilde{\boldsymbol{Z}}_{t_{N}}$ is then denoised for $N$ timesteps to obtain the reconstruction $\boldsymbol{Z}_{t_{0}}$.
+During this process, we store the latent $\boldsymbol{\widetilde{Z}}_{t_{i}}$ and $\boldsymbol{Z}_{t_{i}}$ at each timestep respectively in inversion and denoising processes.
+Then we calculate the Mean Squared Error (MSE) between them.
+The red curve represents the vanilla Rectified Flow inversion and the green curve represents RF-Solver inversion.
+Based on the analysis, we aim to improve inversion accuracy by introducing a more effective sampler, which is more general and fundamental than designing a specific inversion method.
+To this end, we propose RF-Solver.
+Specifically, we note that the exact formulation of the RF ODE solution can be directly derived using the variation of constants method.
 For the nonlinear component of this solution (i.e., the integral of the neural network), we utilize Taylor expansion for estimation.
-By employing higher-order Taylor expansion, the ODE can be solved with reduced error, thereby improving the performance of rectified flow models.
-RF-Solver is a generic sampler that can be seamlessly integrated into any rectified flow model without the need for training or optimization.
-Experimental results demonstrate that RF-Solver not only significantly enhances the accuracy of inversion and reconstruction, but also improves performance on fundamental tasks such as T2I generation.
-Building upon this, we propose RF-Edit to apply RF-Solver in editing tasks.
-Real-world image and video editing require the model to make precise modifications to a source image/video while maintaining its overall structure unchanged.
-This makes editing a more challenging task compared to reconstruction.
-In this scenario, it is inadequate to solely rely on the inverted noises as prior knowledge for editing, which could lead to edited results being excessively influenced by the target prompt, resulting in a completely different output compared with source image/video.
-Addressing this problem, we store the $V$ (value) feature in the self-attention layers at several timesteps during inversion.
-In the process of denoising, these features are used to replace the corresponding features.
-Practically, we design two specific sub-modules for RF-Edit, respectively leveraging FLUX [1] and OpenSora [2] as the backbone for image and video editing.
+By employing higher-order Taylor expansion, the ODE can be solved with reduced error, thereby enhancing the performance of RF models.
+RF-Solver is a generic sampler that can be seamlessly integrated into any rectified flow model without additional training or optimization.
+Experimental results demonstrate that RF-Solver not only significantly enhances the accuracy of inversion and reconstruction (the green curve in Fig.
+2), but also improves performance on fundamental tasks such as T2I generation.
+Building upon this, we propose RF-Edit to leverage RF-Solver in editing tasks.
+Real-world image and video editing require the model to make precise modifications to a source image/video while maintaining its overall structure unchanged, presenting greater challenges than reconstruction.
+In this scenario, it is inadequate to solely rely on the inverted noises as prior knowledge for editing, which could lead to edited results being excessively influenced by the target prompt, diverging significantly from the original source [20, 64].
+Addressing this problem, RF-Edit stores the $\mathcal{V}$ (value) feature in the self-attention layers at several timesteps during inversion.
+These features are used to replace the corresponding features in the denoising process.
+Practically, we design two specific sub-modules for RF-Edit, respectively leveraging the DiT structure of FLUX [2] and OpenSora [3] as the backbones for image and video editing.
 With the effective design of RF-Edit, it demonstrates superior performance in both image and video domains, outperforming various SOTA methods.
 Our core contributions are summarized as follows:
-  • We propose RF-Solver, a training-free sampler that significantly reduces errors in the inversion and reconstruction processes of the rectified-flow model.
-  • We present RF-Edit, which leverages RF-Solver for image and video editing, effectively preserving the structural integrity of the source image/video while achieving high-quality results.
-  • Extensive experiments on images and videos demonstrate the efficacy of our methods, showcasing superior performance in both inversion and high-quality editing compared to various existing baselines.
+  • We propose RF-Solver, a training-free sampler that significantly reduces errors in the inversion and reconstruction processes of rectified-flow models.
+  • We present RF-Edit, which leverages RF-Solver for image and video editing. RF-Edit effectively preserves the structural integrity of the source image/video while delivering high-quality results.
+  • Extensive experiments across a range of tasks demonstrate the efficacy and versatility of our method.
 2 Related Work
@@
 Inversion maps the real visual data, i.e.
-image and video, to a representation in the noise space [14, 50, 35, 44, 34, 23], which is the reverse process of generation.
-This representation captures the essential features and structures of the original data while allowing for flexibility in manipulation for various editing applications.
-Numerous previous inversion approaches have been elaborately designed for diffusion models to achieve remarkable performance.
-The representative method, DDIM inversion [46, 47], adds predicted noise recursively at each forward step, outputting the final state as structured noise.
-To mitigate the discretization error in DDIM inversion, some efforts have been explored from the perspective of optimizing null prompt embeddings [35] and latent variables [44].
-Negative prompt inversion [34] accelerates the inversion process at the cost of fidelity.
-Despite the success of inversion in diffusion models, the exploration of inversion in SOTA rectified flow models like FLUX and OpenSora is limited.
-RF-prior [57] uses the score distillation to invert the image while it requires a number of optimizing steps.
-More recently, [45] introduces an additional vector field that is conditioned on the source image to improve the inversion.
-However, its performance improvement mainly stems from the information leakage of the source image during the inversion and reconstruction process.
-The error from the original vector field of rectified flow still persists, which would limit the performance of such method on various downstream tasks.
+image and video, to representations in noise space, which is the reverse process of generation.
+The representative method, DDIM inversion [60, 61], adds predicted noise recursively at each forward step.
+Many efforts [14, 65, 47, 57, 46, 27, 46, 44, 19] have been made to mitigate the discretization error in DDIM inversion.
+Despite the effectiveness of inversion in diffusion models, the exploration of inversion in SOTA rectified flow models like FLUX and OpenSora is limited.
+RF-prior [76] uses the score distillation to invert the image while it requires many optimizing steps.
+More recently, [58] introduces an additional vector field conditioned on the source image to improve the inversion.
+However, the error from the original vector field of rectified flow still persists, which would limit the performance of such method on various downstream tasks.
+In contrast, we aim to directly mitigate the error from the original vector field in this work.
 2.2 Image and Video Editing
-Training-free methods for image and video editing [21, 48] have gained increasing popularity due to their efficiency and effectiveness.
-These methods usually invert the source image/video into Gaussian noises and then denoises it conditioned on the target prompt.
-Existing image editing methods focus on prompt refinement [41, 52], attention-sharing mechanism [17, 37, 7, 49], mask guidance [5, 11, 20, 28], and noise initialization [6, 58].
-Video editing introduces additional complexities due to the need for maintaining temporal consistency, making it a more challenging task.
+Training-free methods for image and video editing [25, 62] have gained increasing popularity for their efficiency and effectiveness.
+Existing image editing methods focus on prompt refinement [55, 69], attention-sharing mechanism [20, 50, 18, 64], mask guidance [4, 10, 24, 37, 19], and noise initialization [5, 77].
+Video editing introduces additional complexities in maintaining temporal consistency, making it a more challenging task.
 Existing video editing methods focus on attention injection [53, 70, 41], motion guidance [9, 16, 75, 68], latent manipulation [81, 74, 30, 8], and canonical representation [7, 36, 49, 31].
-To date, the editing performance of rectified-flow-based diffusion transformers has remained largely under-explored.
-Although [45] employs FLUX for image editing, its performance is limited to simple tasks such as stylization and face editing, and it struggles to effectively maintain the structural information of source images.
-Moreover, currently there is no research exploring the video editing capabilities of rectified-flow-based models.
+To date, the editing performance of RF-based diffusion transformers has remained largely under-explored.
+Although [58] employs FLUX [2] for image editing, its performance is limited to simple tasks such as stylization and face editing while often failing to effectively maintain the structural information of source images.
+Moreover, currently there is no research exploring the video editing capabilities of RF-based models.
 3 Method
-In this section, we present our methods in detail.
-First, we introduce the proposed RF-Solver, which significantly enhances the precision of inversion and reconstruction.
-Subsequently, we introduce RF-Edit, an extension of RF-Solver designed to enable high-quality image and video editing.
+In this section, we present our method in detail.
+First, we introduce RF-Solver, which significantly enhances the precision of inversion and reconstruction.
+Subsequently, we present RF-Edit, an extension of RF-Solver designed to enable high-quality image and video editing.
 3.1 Preliminaries
-Rectified Flow [32] facilitates the transport between the Gaussian Noises distributions $\pi_{0}$ and real data distribution $\pi_{1}$ along a straight path.
-This is achieved by learning a forward-simulating system given by $\frac{d\boldsymbol{Z}_{t}}{dt}=v(\boldsymbol{Z}_{t},t),t\in[0,1]$ which maps $\boldsymbol{Z}_{0}\in\pi_{0}$ to $\boldsymbol{Z}_{1}\in\pi_{1}$.
-In practice, the velocity field $v$ is parameterized by a neural network $\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t},t)$.
-During training, given $\boldsymbol{X}_{0}\in\pi_{0}$ and $\boldsymbol{X}_{1}\in\pi_{1}$, the forward process of rectified flow is implemented through a simple linear combination:
-$\boldsymbol{X}_{t}=t\boldsymbol{X}_{1}+(1-t)\boldsymbol{X}_{0},t\in[0,1].$ (1)
-The rectified flow ODE can be derived by differentiating Eq. 1 with respect to $t$, as follows:
-$\frac{d\boldsymbol{X}_{t}}{dt}=\boldsymbol{X}_{1}-\boldsymbol{X}_{0},t\in[0,1].$ (2)
-Here, $\boldsymbol{X}_{1}-\boldsymbol{X}_{0}$ serves as the ground truth, and the network is optimized through mean square error:
+Rectified Flow (RF) [43] facilitates the transition between the real data distribution $\pi_{0}$ and Gaussian Noises distribution $\pi_{1}$ along a straight path.
+This is achieved by learning a forward-simulating system defined by the ODE: $d\boldsymbol{Z}_{t}=v(\boldsymbol{Z}_{t},t)dt,t\in[0,1]$, which maps $\boldsymbol{Z}_{1}\in\pi_{1}$ to $\boldsymbol{Z}_{0}\in\pi_{0}$.
+In practice, the velocity field $v$ is parameterized by a neural network $\boldsymbol{v}_{\theta}$.
+During training, given empirical observations of two distributions $\boldsymbol{X}_{0}\sim\pi_{0}$, $\boldsymbol{X}_{1}\sim\pi_{1}$ and $t\in[0,1]$, the forward process (i.e.
+adding noise) of rectified flow is defined by a simple linear combination: $\boldsymbol{X}_{t}=t\boldsymbol{X}_{1}+(1-t)\boldsymbol{X}_{0}$.
+The differential form of the equation is given by: $d\boldsymbol{X}_{t}=(\boldsymbol{X}_{1}-\boldsymbol{X}_{0})dt$.
+Consequently, the training process optimizes the network by solving the least squares regression problem, which fits the $\boldsymbol{v}_{\theta}$ with $(\boldsymbol{X}_{1}-\boldsymbol{X}_{0})$:
 […]
-During the sampling process, the ODE is discretized and solved using Euler method.
-Specifically, the rectified flow model starts with Gaussian noises $\boldsymbol{Z}_{t_{N}}\in\mathcal{N}(0,\boldsymbol{I})$.
-Given a series of discrete $N$ timesteps $t=\{t_{N},...,t_{0}\}$, the model iteratively predicts $\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}}),i\in\{N,\cdots,1\}$ and then takes a step forward until generating the images $\boldsymbol{Z}_{t_{0}}$, with the following recurrence relation:
+In the sampling process, the ODE is discretized and solved using the Euler method.
+Specifically, the rectified flow model starts with a Gaussian noise sample $\boldsymbol{Z}_{t_{N}}\in\mathcal{N}(0,\boldsymbol{I})$.
+Given a series of $N$ discrete timesteps $t=\{t_{N},...,t_{0}\}$, the model iteratively predicts $\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$ for $i\in\{N,\cdots,1\}$ and then takes a step forward until generating the images $\boldsymbol{Z}_{t_{0}}$, with the following recurrence relation:
 $\boldsymbol{Z}_{t_{i-1}}=\boldsymbol{Z}_{t_{i}}+(t_{i-1}-t_{i})\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i}).$ (2)
-The definition and training process of Rectified Flow is designed to establish a nearly linear transition trajectory between two distributions, enabling efficient generation with significantly fewer steps compared to DDPMs [19].
+The RF model can generate high-quality images in much fewer timesteps compared to DDPM [22], owing to the nearly linear transition trajectory established during training.
+With these effective designs, RF model illustrates great potential in the field of T2I and T2V generation [2, 3].
 3.2 RF-Solver
-Figure 2: Analysis of the Inversion-Reconstruction Process.
-Inversion begins with the source image latent $\widetilde{\boldsymbol{{Z}}}_{t_{0}}$ as the input and progressively add noise, obtaining $\widetilde{\boldsymbol{{Z}}}_{t_{N}}\in\mathcal{N}(0,\boldsymbol{I})$.
-$\widetilde{\boldsymbol{Z}}_{t_{N}}$ is then denoised for $N$ timesteps to obtain the reconstruction $\boldsymbol{Z}_{t_{0}}$.
-During this process, we store the latent $\boldsymbol{\widetilde{Z}}_{t_{i}}$ and $\boldsymbol{Z}_{t_{i}}$ at each time step and calculate the mean squared error between them.
-Figure 3: RF-Edit Pipelines.
-The RF-Edit framework comprises two sub-modules, respectively applied to FLUX [1] for image editing and OpenSora [2] for video editing.
-For image editing, we share the feature of $\widetilde{\mathcal{V}}$ within the single block self-attention of the FLUX [1] backbone.
-For video editing, we share the feature of $\widetilde{\mathcal{V}}$ in the spatial attention of the OpenSora [2] backbone.
-The vanilla rectified flow (RF) sampler demonstrates strong performance in image and video generation.
+The vanilla RF sampler demonstrates strong performance in image and video generation.
 However, when applied to inversion and reconstruction tasks, we observe significant error accumulation at each timestep.
 This results in reconstructions that diverge notably from the original image (see Fig.
-2), further limiting the performance of RF-based models in various downstream tasks, such as image and video editing.
-Delving into this problem, we notice that the inversion and reconstruction processes in rectified flow rely on estimating an approximate solution of the rectified flow ODE at each timestep (see Eq.
-4).
-Obtaining more precise solutions for the ODE would effectively mitigate these errors, leading to improved reconstruction quality.
-Based on this analysis, we start by carefully examining the differential form of the Rectified flow: $\frac{d\boldsymbol{Z}_{t}}{dt}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t},t)$.
+2).
+This severely limits the performance of RF models in various inversion-based downstream tasks [20, 66].
+Delving into this problem, we identify that the errors stem from the process of estimating the approximate solution for the rectified flow ODE [67, 40], which is formulated by Eq.
+2 in existing methods [2, 3].
+Consequently, obtaining more precise solutions for the ODE would effectively mitigate these errors, leading to improved reconstruction quality.
+Based on this analysis, we start by carefully examining the differential form of the Rectified flow: ${d\boldsymbol{Z}_{t}}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t},t)dt$.
 This ODE is discretized in the sampling process.
@@
 Through the above process, the network prediction term and its higher-order derivatives are separated from the integral.
-Then we notice that the remaining portion in the integral can be computed analytically:
+Then we notice that the remaining component in the integral can be computed analytically:
 […]
@@
 3.2 indicates that to estimate $\boldsymbol{Z}_{t_{i-1}}$, we need to obtain the $k$-th order derivatives $\{\boldsymbol{v}^{(k)}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})\}$ for $k\in\{0,\cdots,n-1\}$.
+When $n=1$, the formula reduces to the standard rectified flow (i.e., Eq. 2). In our experiments, we find that setting $n=2$ effectively mitigates the errors, yielding:
+$\boldsymbol{Z}_{t_{i-1}}=\boldsymbol{Z}_{t_{i}}$ $+({t_{i-1}}-{t_{i}})\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i})$
+$+\frac{1}{2}(t_{i-1}-t_{i})^{2}\boldsymbol{v}_{\theta}^{(1)}(\boldsymbol{Z}_{t_{i}},t_{i}).$ (8)
+Figure 3: RF-Edit pipelines for image editing and video editing.
+We design two sub-modules for applying RF-Edit to (a).
+Image editing with FLUX [2] and (b).
+Video editing with OpenSora [3].
+Note that for FLUX, there are multiple Double Blocks, followed by multiple Single Blocks.
+For OpenSora, there are multiple OpenSora DiT blocks.
+For simplicity, only one block of each type is depicted in the figure.
+Note that $\boldsymbol{v}_{\theta}^{(1)}$ in Sec.
+3.2 is the first-order derivative of the network prediction term $\boldsymbol{v}_{\theta}$, which cannot be analytically derived due to the complex architecture of the neural network.
+To estimate this term, we first obtain the network prediction $\hat{\boldsymbol{v}}_{t_{i}}$ at the timestep $t_{i}$, i.e., $\hat{\boldsymbol{v}}_{t_{i}}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$.
+Then we step forward a small timestep $\Delta t=\frac{1}{2}(t_{i-1}-t_{i})$, and update the latents to obtain $\boldsymbol{Z}_{{t_{i}}+\Delta t}=\boldsymbol{Z}_{t_{i}}+\Delta t\cdot\hat{\boldsymbol{v}}_{t_{i}}$.
+Subsequently, we calculate an additional prediction of the network at the timestep $t_{i}+\Delta t$, i.e., $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}+\Delta t},{t_{i}+\Delta t})$.
+With $\hat{\boldsymbol{v}}_{t_{i}}$ and $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}$, the first-order derivative of ${\boldsymbol{v}}_{\theta}$ at the timestep $t_{i}$ can be estimated as: $\boldsymbol{v}_{\theta}^{(1)}(\boldsymbol{Z}_{t_{i}},t_{i})=\frac{\hat{\boldsymbol{v}}_{t_{i}+\Delta t}-\hat{\boldsymbol{v}}_{t_{i}}}{\Delta t}$.
+Substituting this formulation into Sec.
+3.2 results in the practical implementation of the RF-Solver algorithm.
+The complete sampling process for RF-Solver is presented in Algorithm 1.
 Algorithm 1 Sampling process of RF-Solver
@@
   $\boldsymbol{\hat{v}}_{t_{i}+\Delta t_{i}}\leftarrow\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}+\Delta t_{i}},t_{i}+\Delta t_{i})$
-  $\boldsymbol{v}_{t_{i}}^{(1)}\leftarrow({\boldsymbol{\hat{v}}_{t_{i}}-\boldsymbol{\hat{v}}_{t_{i}+\Delta t_{i}}})/{\Delta t_{i}}$triangleright Calculating the Derivatives
+  $\boldsymbol{v}_{t_{i}}^{(1)}\leftarrow({\boldsymbol{\hat{v}}_{t_{i}+\Delta t_{i}}-\boldsymbol{\hat{v}}_{t_{i}}})/{\Delta t_{i}}$triangleright Calculating the Derivatives
   $\boldsymbol{Z}_{t_{i-1}}\leftarrow\boldsymbol{Z}_{t_{i}}+(t_{i-1}-t_{i})\boldsymbol{\hat{v}}_{t_{i}}+\frac{1}{2}(t_{i-1}-t_{i})^{2}\boldsymbol{v}_{t_{i}}^{(1)}$
 Output: $\boldsymbol{Z}_{0}$
-When $n=1$, the formula reduces to the standard rectified flow (Eq. 4). In our experiments, we find that setting $n=2$ effectively mitigates the errors, yielding the following formula:
-$\boldsymbol{Z}_{t_{i-1}}=\boldsymbol{Z}_{t_{i}}$ $+({t_{i-1}}-{t_{i}})\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i})$
-$+\frac{1}{2}(t_{i-1}-t_{i})^{2}\boldsymbol{v}_{\theta}^{(1)}(\boldsymbol{Z}_{t_{i}},t_{i}).$ (10)
-Note that $\boldsymbol{v}_{\theta}^{(1)}$ is the first-order derivative of the network prediction term $\boldsymbol{v}_{\theta}$, which cannot be analytically derived due to the complex architecture of the neural network.
-To estimate this term, we first obtain the network prediction $\hat{\boldsymbol{v}}_{t_{i}}$ at the timestep $t_{i}$, i.e., $\hat{\boldsymbol{v}}_{t_{i}}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$.
-Then we step forward a small timestep $\Delta t=\frac{1}{2}(t_{i-1}-t_{i})$, and update the latents to obtain $\boldsymbol{Z}_{{t_{i}}+\Delta t}=\boldsymbol{Z}_{t_{i}}+\Delta t\cdot\hat{\boldsymbol{v}}_{t_{i}}$.
-Next, we calculate an additional prediction of the network at the timestep $t_{i}+\Delta t$, i.e., $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}+\Delta t},{t_{i}+\Delta t})$.
-With $\hat{\boldsymbol{v}}_{t_{i}}$ and $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}$, the first-order derivative of ${\boldsymbol{v}}_{\theta}$ at the timestep $t_{i}$ can be estimated as:
-$\boldsymbol{v}_{\theta}^{(1)}(\boldsymbol{Z}_{t_{i}},t_{i})=\frac{\hat{\boldsymbol{v}}_{t_{i}+\Delta t}-\hat{\boldsymbol{v}}_{t_{i}}}{\Delta t}.$ (11)
-Substituting Eq. 11 into Sec. 3.2 results in the practical implementation of the RF-Solver algorithm. The complete sampling process for RF-Solver is presented in Algorithm 1.
-Besides sampling, inversion seeks to map data back into noise, which reverses the sampling process.
+Obtaining the sampling form of RF-Solver, we further derive its inversion form.
+Inversion maps data back into noise, which reverses the sampling process.
 Following previous methods for DDIM inversion [60, 12], the ODE process can be directly reversed in the limit of small steps.
 Based on this assumption, the inversion process of RF-Solver (Sec.
-3.2) can be directly transformed as:
+3.2) can be directly derived as:
 $\widetilde{\boldsymbol{Z}}_{t_{i+1}}=\widetilde{\boldsymbol{Z}}_{t_{i}}$ $+({t_{i+1}}-{t_{i}})\boldsymbol{v}_{\theta}(\widetilde{\boldsymbol{Z}}_{t_{i}},t_{i})$
@@
 where $\widetilde{\boldsymbol{Z}}_{t_{i}}$ and $\widetilde{\boldsymbol{Z}}_{t_{i+1}}$ denotes the latents during inversion.
-Through this high order expansion, the error of the ODE solution in each timestep is reduced from $\mathcal{O}\big{(}(h_{i})^{2}\big{)}$ to $\mathcal{O}\big{(}(h_{i})^{3}\big{)}$, leading to improved performance, particularly in inversion and reconstruction (see Fig.
+Through the high order expansion, the error of the ODE solution in each timestep is reduced from $\mathcal{O}\big{(}(h_{i})^{2}\big{)}$ to $\mathcal{O}\big{(}(h_{i})^{3}\big{)}$, leading to improved performance, particularly in inversion and reconstruction (see Fig.
 2).
@@
 3.3 RF-Edit
-Incorporating higher-order terms enables RF-Solver to significantly reduce errors in the ODE-solving process, improving both sampling quality and inversion accuracy.
-Furthermore, we extend the application of RF-Solver to real-world image and video editing tasks, which are more challenging than reconstruction.
-In these scenarios, maintaining the content and structure of the original image is crucial.
-For instance, when adding new objects to a source image, other unrelated regions should remain unaffected by the editing process.
-However, directly applying RF-Solver during the inversion and denoising stages may cause the model to be overly influenced by the target prompt, leading to unintended modifications in other parts of the image or video (e.g., altering the unrelated which is not mentioned in the editing prompt).
-This issue is common across various existing editing methods [45, 17, 49].
-To address this problem, we propose RF-Edit, which incorporates features from the inversion process into the denoising procedure.
-Specifically, during the last $n$ steps of inversion, we extract and store the Value feature $\{\mathcal{{\widetilde{V}}}_{t_{k}}^{m}\}$ and $\{\mathcal{{\widetilde{V}}}_{t_{k}+\Delta t_{k}}^{m}\}$ from the self-attention layers in the last $M$ transformer blocks at each timestep $k$.
-Here, $k\in\{n,n+1,\cdots,N\}$ and $m\in\{1,\cdots,M\}$.
-This process can be formulated as follows:
+Incorporating higher-order terms enables RF-Solver to significantly reduce errors in the ODE-solving process, thereby enhancing both sampling quality and inversion accuracy.
+Furthermore, we extend the application of RF-Solver to the more complex real-world image and video editing tasks, which present greater challenges than reconstruction.
+In such scenarios, preserving the content and structure of the original image is crucial.
+For example, when replacing an object in a source image with another one, regions unrelated to the object in this image are expected to remain unaffected by the editing process.
+However, directly applying RF-Solver during the inversion and denoising stages may cause the model to be overly influenced by the target prompt, resulting in unintended modifications in other regions of the source image or video.
+Similar issues are common across various existing editing methods [58, 20, 64].
+To address this problem, we propose RF-Edit, which builds upon the diffusion transformer architecture.
+Specifically, we focus on the self-attention layer in the last $M$ transformer blocks of $\boldsymbol{v}_{\theta}$ at the last $n$ timesteps during inversion.
+The self-attention operation can be formulated by:
+$\boldsymbol{\widetilde{F}}^{m}_{t_{k}}=\text{Attention}(\mathcal{{\widetilde{Q}}}_{t_{k}}^{m},\mathcal{{\widetilde{K}}}_{t_{k}}^{m},\mathcal{{\widetilde{V}}}_{t_{k}}^{m}).$ (10)
+Here, $k\in\{N-n,\cdots,N\}$, and $m\in\{1,\cdots,M\}$, $\boldsymbol{\widetilde{F}}^{m}_{t_{k}}$ denotes the output feature of the self-attention module and $\mathcal{{\widetilde{Q}}}_{t_{k}}^{m},\mathcal{{\widetilde{K}}}_{t_{k}}^{m},\mathcal{{\widetilde{V}}}_{t_{k}}^{m}$ represent query, key and value for attention during the inversion process, respectively.
+We extract and store the Value feature $\{\mathcal{{\widetilde{V}}}_{t_{k}}^{m}\}$ and $\{\mathcal{{\widetilde{V}}}_{t_{k}+\Delta t_{k}}^{m}\}$ in the process of RF-Solver algorithm (Algorithm 1):
 $\{\mathcal{{\widetilde{V}}}_{t_{k}}^{m}\}$ $=\text{Extract}\big{(}\boldsymbol{v}_{\theta}(\boldsymbol{\widetilde{Z}}_{t_{k}},t_{k})\big{)}$ (11)
-$\{\mathcal{{\widetilde{V}}}^{m}_{t_{k}+\Delta t_{k}}\}$ $=\text{Extract}\big{(}\boldsymbol{v}_{\theta}(\boldsymbol{\widetilde{Z}}_{t_{k}+\Delta t_{k}},t_{k}+\Delta t_{k})\big{)},$ (14)
-where $\boldsymbol{\widetilde{Z}}_{t_{k}}$ and $\boldsymbol{\widetilde{Z}}_{t_{k}+\Delta t_{k}}$ denote the latents during inversion for RF-Solver.
-During the first $n$ timesteps of denoising, considering the $m$th transformer block at the timestep $k$, the original self-attention in the network $\boldsymbol{v}_{\theta}$ can be formulated as:
+$\{\mathcal{{\widetilde{V}}}^{m}_{t_{k}+\Delta t_{k}}\}$ $=\text{Extract}\big{(}\boldsymbol{v}_{\theta}(\boldsymbol{\widetilde{Z}}_{t_{k}+\Delta t_{k}},t_{k}+\Delta t_{k})\big{)}.$ (12)
+During the first $n$ timesteps of denoising, considering the $m$th transformer block at the timestep $k$, the original self-attention can be formulated as:
 $\boldsymbol{F}^{m}_{t_{k}}=\text{Attention}(\mathcal{{Q}}_{t_{k}}^{m},\mathcal{{K}}_{t_{k}}^{m},\mathcal{{V}}_{t_{k}}^{m}),$ (13)
 where $\boldsymbol{F}^{m}_{t_{k}}$ denotes the output feature of the self-attention module and $\mathcal{{Q}}_{t_{k}}^{m},\mathcal{{K}}_{t_{k}}^{m},\mathcal{{V}}_{t_{k}}^{m}$ represent query, key and value for attention during the denoising process, respectively.
-Figure 4: Qualitative Results of Image and Video Reconstruction.
-Our method (The second row) illustrates better performance compared with the rectified flow baselines (The third row) on both image and video reconstruction.
+Figure 4: Qualitative results of image and video reconstruction. Our method (the second row) demonstrates superior performance compared to the vanilla rectified flow baselines (the third row).
 In RF-Edit, the above self-attention mechanism is modified to cross-attention where $\mathcal{{V}}_{t_{k}}^{m}$ is replaced by $\mathcal{{\widetilde{V}}}_{t_{k}}^{m}$,
@@
 ${{\boldsymbol{F}}^{m\prime}_{t_{k+\Delta t_{k}}}}=\text{Attention}(\mathcal{{Q}}_{t_{k+\Delta t_{k}}}^{m},\mathcal{{K}}_{{k+\Delta t_{k}}}^{m},\mathcal{{\widetilde{V}}}_{{k+\Delta t_{k}}}^{m}).$ (15)
-The proposed RF-Edit framework enables high-quality editing while preserving structural information.
-Building on this concept, we design two sub-modules for RF-Edit, specifically tailored for image editing and video editing.
-For image editing, we use FLUX [1] as the backbone, which comprises several double blocks and single blocks.
+The proposed RF-Edit framework enables high-quality editing while effectively preserving the structural information of the source image/video.
+Building on this concept, we design two sub-modules for RF-Edit, specifically tailored for image editing and video editing (Fig.
+3).
+For image editing, RF-Edit employs FLUX [2] as the backbone, which comprises several double blocks and single blocks.
 Double blocks independently modulate text and image features, while single blocks concatenate these features for unified modulation.
 In this architecture, RF-Edit shares features within the single blocks, as they capture information from both the source image and the source prompt, enhancing the ability of the model to preserve the structural information of the source image.
-For video editing, we employ OpenSora [2] as the backbone.
+For video editing, RF-Edit employs OpenSora [3] as the backbone.
 The DiT blocks in OpenSora include spatial attention, temporal attention, and text cross-attention.
@@
 Baselines.
-We select the vanilla Rectified Flow sampler as the primary baseline for all tasks.
-Additionally, for image editing, we compare our method with P2P [17], DiffEdit [12], SDEdit [33], PnP [49], Pix2pix [37] and RF-Inversion [45].
-For video editing tasks, we compare our methods with FateZero [39], FLATTEN [10], COVE [51], RAVE [25], Tokenflow [15].
-Detailed experimental settings of these methods are provided in the Appendix.
-Figure 5: Qualitative Results of Text-to-Image Generation. With the RF-Solver, the model can generate images with higher quality (The first row) than baselines (The Second row).
-Metric FID ($\downarrow$) Clip Score ($\uparrow$)
-RF 26.55 31.01
-Ours 25.45 31.09
-Table 1: Quantitive results on Text-to-Image Generation. RF-Solver outperforms the vanilla RF sampler.
-Figure 6: Qualititive Comparison of Image Editing.
-With RF-Solver and feature-sharing mechanism in RF-Edit, our method can successfully handle various kinds of image editing cases, outperforming the previous SOTA methods.
-Zoom-in for the best views.
+We use the vanilla Rectified Flow sampler (Euler Sampler) as the primary baseline for all tasks.
+Besides, for sampling and inversion, we compare our method with DPM-Solver++ [44] and the Heun sampler (a 2-order ODE solver).
+For image editing, we compare our method with P2P [20], DiffEdit [11], SDEdit [45], PnP [64], Pix2pix [50] and RF-Inversion [58].
+For video editing tasks, we compare our method with FateZero [53], FLATTEN [9], COVE [68], RAVE [30], Tokenflow [16].
+Detailed experimental settings of baselines are provided in the Appendix.
+Figure 5: Qualitative results of text-to-image generation. By employing the RF-Solver, the model is able to generate images with higher quality (the first row) than baselines (the second row).
 Implementation Details.
 In the experiment, we adopt the guidance-distilled variant of FLUX [2] for image tasks and OpenSora [3] for video tasks.
-For sampling, we set the denoising step to 10 for our method.
-For reconstruction and editing, the inversion and denoising steps are set to 30.
 The derivative computation in RF-Solver requires an additional forward pass, resulting in the network needing to forward twice at each timestep.
-As a result, when comparing our method with the Rectified Flow baselines, we set the number of timestep for the vanilla Rectified Flow to be twice that of our method in order to maintain the same number of function evaluations (NFE) to ensure a fair comparison.
-Both the baseline and our method are conducted on a single A100 GPU with 40GB memory.
+As a result, when comparing our method with the Rectified Flow baselines, we set the number of timesteps for the vanilla Rectified Flow to be twice that of our method to ensure a fair comparison under the same number of function evaluations (NFE).
+More information is provided in the Appendix.
 Evaluation Metrics For text-to-image sampling, we randomly select 10k images from the MSCOCO validation set [39] and report the FID [21] and CLIP Score [54].
 For the inversion and reconstruction task, we report the Mean Square Error (MSE), LPIPS [80], SSIM [72], and PSNR [28].
-For image editing tasks, we report the Clip Score [40], which reflects whether the edited images align with the target prompt, and LPIPS [60], which reflects whether the edited images preserve the content of the source images.
+For image editing tasks, we report the CLIP Score [54] and LPIPS [80].
 For video editing tasks, we adopt the metric proposed by [26], including Subject Consistency (SC), Motion Smoothness (MS), Aesthetic Quality (AQ), and Imaging Quality (IQ).
-SC and MS assess the temporal consistency of the edited video, while AQ and IQ assess the visual quality.
+Detailed explanations of these metrics are provided in the Appendix.
 4.2 Text-to-Image Sampling
-We compare the performance of our method with the vanilla rectified flow on the text-to-image generation task.
+We compare the performance of our method with DPMSolver++ [44], the vanilla RF sampler, and Heun sampler on the text-to-image generation task.
 Both the quantitative (Tab.
@@
 5) demonstrate the superior performance of RF-Solver in fundamental T2I generation tasks, producing higher-quality images that align more closely with human cognition.
+Figure 6: Qualitative comparison of image editing.
+With RF-Solver and feature-sharing mechanism in RF-Edit, our method can successfully handle various kinds of image editing cases, outperforming the previous SOTA methods.
+Zoom in for the best views.
+DPMSolver++ RF RF-Heun Ours
+FID ($\downarrow$) 24.63 25.33 24.40 24.03
+CLIP Score ($\uparrow$) 30.62 31.01 31.03 31.09
+Table 1: Quantitative results on text-to-image generation. RF-Solver outperforms several baselines.
 4.3 Inversion and Reconstruction
-We conduct experiments on inversion and reconstruction for both image and video modalities, comparing our methods with the vanilla rectified flow sampler.
-For image inversion and reconstruction, we use images from the MSCOCO dataset, and for video inversion and reconstruction, we select videos from social media platforms such as TikTok and other publicly available sources [3, 4].
-We use GPT-4.0 to annotate the content of the images and videos in detail, and then manually polish the GPT-generated content.
-These annotations are used as the source and target prompts for inversion and reconstruction.
-Quantitive Comparison.
+We conduct experiments on inversion and reconstruction for both image and video modalities, comparing our method with the vanilla RF sampler and the Heun sampler.
+Quantitative Comparison.
 The quantitative comparisons (Tab.
-2) are conducted on both images and videos to illustrate the similarity between the source and reconstruction results.
-Our method demonstrates superior performance across all four metrics compared with vanilla rectified flow.
+2) are conducted to reflect the similarity between the source and reconstruction results.
+Our method demonstrates superior performance across all four metrics compared with the vanilla RF sampler and Heun sampler.
 Mehtod MSE ($\downarrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$)
 image RF 0.0268 0.6253 0.7626 28.28
+RF-Heun 0.0117 0.4696 0.8924 29.67
 Ours 0.0094 0.4242 0.9271 29.83
 video RF 0.0206 0.4159 0.8134 18.12
+RF-Heun 0.0156 0.3554 0.8711 18.29
 Ours 0.0139 0.3299 0.8805 18.32
-Table 2: Quantitative Results on Inversion and Reconstruction. Our methods significantly improve the reconstruction accuracy of both images and videos.
+Table 2: Quantitative results on inversion and reconstruction. Our method significantly improves the accuracy of reconstruction for both images and videos.
 Qualitative Comparison.
@@
 As illustrated in Fig.
-4(a), the image reconstruction results using vanilla rectified flow exhibit noticeable drift from the source image, with significant alterations to the appearance of subjects such as cake, church, train, and jeep.
+4(a), the image reconstruction results using vanilla rectified flow exhibit noticeable drift from the source image, with significant alterations to the appearance of subjects in the image.
 For video reconstruction, as shown in Fig.
 4(b), the baseline reconstruction results suffer from distortion.
-In contrast, RF-Solver significantly alleviate these issues, achieving more satisfactory results.
-Figure 7: Qualititive Comparison of Video Editing.
-The first video has 200 frames with a resolution of $512\times 512$, and the second video has 60 frames with a resolution of $1024\times 768$ (we compress the frames for a neat layout).
-Our method outperforms various previous SOTA video editing methods, especially excelling in dealing with complex prompts.
+In contrast, RF-Solver significantly alleviates these issues, achieving more satisfactory results.
+Figure 7: Qualitative comparison of video editing.
+The first video comprises 200 frames with a resolution of $512\times 512$, while the second video contains 60 frames with a resolution of $1024\times 768$ (frames are compressed for a neat layout in the figure).
 4.4 Editing
-We conduct experiments to evaluate the image and video editing performance of our methods.
+We conduct experiments to evaluate the image and video editing performance of our method.
 Image editing usually involves replacing the subject in the image with another one, adding new items, and global editing.
@@
 For global editing such as style transfer, the overall structure of the source image is expected to remain unchanged.
-Compared to image editing, video editing is much more challenging due to the complexity of modeling temporal motions in videos.
-Recent mainstream video editing methods usually focus on replacing the subjects and performing global editing.
+Recent mainstream video editing methods usually focus on replacing the subjects and performing global editing for the source video.
 Quantitative Comparison.
-For image editing, we perform quantitative comparison between our methods and baselines (Tab.
-3), reporting the Clip score and LPIPS.
-Our method outperforms all other methods in Clip score, indicating that the edited images align well with the user-provided prompts.
+In image editing, Our method outperforms all other methods in CLIP score (Tab.
+3), indicating that the edited images align well with the user-provided prompts.
 For LPIPS, it is noted that PnP [64] has a much lower value than all other methods.
@@
 6), it can be seen that PnP is only suitable for editing cases that do not significantly modify the structure or shape of the source image (such as changing red roses into yellow sunflowers).
-For shape editing, such as modifying a car into a motorbike, PnP fails to edit this case, resulting in an image very similar to the source.
+It fails in the case of shape editing, resulting in an image very similar to the source.
 Consequently, although PnP has the lowest LPIPS score, its CLIP score is the lowest.
@@
 CLIP Score ($\uparrow$) 30.70 32.68 31.61 30.58 32.33 33.02 33.66
-Table 3: Quantitative Results of Image Editing. RF-Edit can effectively edit the images according to the prompts while keeping the unrelated regions unchanged.
-For video editing, we compare our methods with baseline methods using the VBench [22] metrics (Tab.
+Table 3: Quantitative results of image editing. RF-Edit effectively edit the images according to the prompts while preserving the integrity of unrelated regions.
+For video editing, RF-Edit achieves higher scores on the popular VBench [26] metrics (Tab.
 4).
-The results illustrate that our methods successfully maintain temporal consistency in long videos, achieving the highest Subject Consistency (SC) and Motion Smoothness (MS) scores.
-Additionally, our method demonstrates superior visual quality, outperforming the baselines at Aesthetic Quality (AQ) and Image Quality (IQ).
+The results illustrate that our method successfully maintains temporal consistency while demonstrating superior visual quality.
 FateZero Flatten COVE RAVE Tokenflow Ours
@@
 IQ ($\uparrow$) 0.6898 0.7024 0.7163 0.6917 0.7128 0.7207
-Table 4: Quantitative Results of Video Editing. RF-Edit outperforms a number of previous SOTA video editing methods.
+Table 4: Quantitative results of video editing. RF-Edit outperforms several previous SOTA video editing methods.
 Qualitative Comparison.
-For image editing, we compare the performance of our methods with several baselines across different types of editing tasks (Fig.
+For image editing, we compare the performance of our method with several baselines across different types of editing tasks (Fig.
 6).
 The baseline methods often suffer from background changes or fail to perform the desired edits.
-In contrast, our methods demonstrate satisfying performance, effectively achieves a balanced trade-off between the fidelity to the target prompt and preservation of the source image.
-To be noticed, although RF-inversion [45] also uses the rectified flow model for image editing (third row in Fig.
-6), the structure of the source image which is unrelated to editing prompt (such as background and human appearance) is modified obviously.
-For video editing, we primarily evaluate the performance of our methods on long videos (200 frames) and high-resolution videos ($1280\times 768$).
-Furthermore, we assess the performance on complicated videos and prompts where there are multiple objects in the video, and the user has different editing requirements for each object.
+In contrast, our method demonstrates satisfying performance, effectively achieving a balanced trade-off between the fidelity to the target prompt and the preservation of the source image.
+For video editing, we primarily evaluate the performance of our method on long videos (200 frames) and high-resolution videos ($1280\times 768$).
 The qualitative results are shown in Fig.
 7.
-Our method successfully handles complicated editing cases (e.g., modifying the leftmost lion among three lions into a white polar bear and changing the other two small lions into orange tiger cubs), whereas all other baseline methods fail in this scenario.
-Our method also demonstrates strong performance in global editing tasks, such as transforming scenes into autumn.
+RF-Edit illustrates impressive performance in handling complicated editing cases (e.g., modifying the leftmost lion among three lions into a white polar bear and changing the other two small lions into orange tiger cubs), whereas all other baseline methods fail in this scenario.
+RF-Edit also demonstrates strong performance in global editing tasks, such as transforming scenes into autumn.
+Metric RF RF-Solver-2 RF-Solver-3
+Sampling FID ($\downarrow$) 25.33 24.03 23.96
+CLIP Score ($\uparrow$) 31.01 31.09 31.09
+Inversion MSE ($\downarrow$) 0.0268 0.0094 0.0131
+LPIPS ($\downarrow$) 0.6253 0.4242 0.4817
+Editing LPIPS ($\downarrow$) 0.1524 0.1494 0.1503
+CLIP Score ($\uparrow$) 32.97 33.66 33.18
+Table 5: Ablation study on the Taylor Expansion order. We select the 2-order expansion (i.e. RF-Solver-2) for various downstream tasks due to its effectiveness and simplicity.
 4.5 Ablation Study
@@
 The second-order expansion demonstrated a significant improvement across various tasks compared to the first-order expansion (i.e., the vanilla rectified flow).
-However, higher-order expansions did not yield further enhancements.
+However, higher-order expansions do not yield further enhancements.
 We speculate that this is primarily due to higher-order Taylor expansions requiring more inference steps per timestep.
 With a fixed NFE, this results in a reduced overall number of timesteps compared to lower-order expansions, leading to suboptimal performance.
-Moreover, computing the higher-order derivatives of $v_{\theta}{t_{i}}$ substantially increases the complexity of the algorithm, posing challenges for practical applications.
+Moreover, computing the higher-order derivatives of $\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$ substantially increases the complexity of the algorithm, posing challenges for practical applications.
 Consequently, we predominantly employed second-order expansion in our experiments.
-Metric RF RF-Solver-2 RF-Solver-3
-Sampling FID ($\downarrow$) 26.55 25.45 25.37
-Clip Score ($\uparrow$) 31.01 31.09 31.09
-Inversion MSE ($\downarrow$) 0.0268 0.0094 0.0131
-LPIPS ($\downarrow$) 0.6253 0.4242 0.4817
-Editing LPIPS ($\downarrow$) 0.1524 0.1494 0.1503
-Clip Score ($\uparrow$) 32.97 33.66 33.18
-Table 5: Ablation Study on the Taylor Expansion Order. We choose the 2-order expansion (i.e. RF-Solver-2) for the downstream tasks for its effectiveness and simplicity.
 Figure 8: Ablation study of feature-sharing step in RF-Edit.
-A too-small feature-sharing step leads to inconsistency between source and target images.
-On the other hand, a too-large feature-sharing step can cause the failure of editing.
+A too-small feature-sharing step results in the inconsistency between source and target images.
+Conversely, a too-large feature-sharing step can lead to the failure of editing.
 Feature Sharing Steps of RF-Edit.
@@
 Based on RF-Solver, we further propose RF-Edit, which achieves high-quality editing performance while effectively preserving the structural information in source images or videos.
-Extensive experiments demonstrate the versatility and effectiveness of our methods.
+Extensive experiments demonstrate the versatility and effectiveness of our method.
+Supplementary Material
+6 Experimental Settings
+6.1 Baselines and Implementation Details
+Text-to-Image Generation.
+We compare our methods with the following baselines: FLUX with the vanilla sampler, Heun Solver, and DPM-Solver.
+The Heun Solver is a second-order ODE solver that can be applied to pretrained rectified flow to solve the ODE more precisely.
+DPM-Solver is a high-order sampler for diffusion ODE, which is not suitable for RF-based models like FLUX.
+As an alternative, we apply the DPM-Solver on Stable Diffusion to evaluate its performance.
+For FLUX with the vanilla sampler and the Heun Solver, we randomly select 10000 images from the MS-COCO validation dataset and use their caption as the prompt for generation.
+The resolution of generated images is $1024\times 1024$.
+For DPM-Solver, we adopt the implementation from the diffuser, adopting its default setting to generate images.
+The total NFE for generating one image is set to 10 for both our method and baselines.
+Inversion.
+We compare the performance of our methods among RF with the vanilla sampler and the Heun sampler.
+For image inversion, we also use the images from the MS-COCO validation set.
+For video inversion, we select videos from social media platforms such as TikTok and other publicly available sources.
+We have observed the quality of the text prompts significantly influence the quality of inversion.
+Consequently, we employ GPT-4o to generate detailed captions for both images and videos, which are then used in the inversion tasks.
+The total NFE for generating one image/video is set to 50 for both our method and baselines.
+Editing.
+For image editing, we compare our methods with RF-inversion and several diffusion-based editing methods.
+For RF-inversion, we adopt the implementation in ComfyUI [1].
+For other baselines, we use their implementation from diffusers.
+For each baseline, we adjust the relevant hyper-parameters to achieve optimal results.
+For video editing, we use the official codes of all the baseline methods and tune the hyper-parameters to achieve satisfactory results.
+For image editing, we share the features of the last 19 single blocks in FLUX.
+For video editing, we share the features of the last 14 blocks in Open-Sora.
+We adjust the hyper-parameter of feature-sharing steps to achieve better results for both image and video editing.
+Figure 9: Stylization Results.
+6.2 Evaluation Metrics
+For text-to-image sampling, we report Fréchet Inception Distance (FID) and CLIP Scores.
+The FID is a metric used to evaluate the quality of generated images by assessing the similarity between the distributions of real and generated image features, typically extracted using a pre-trained Inception network.
+The CLIP Score evaluates the alignment between generated images and textual descriptions by measuring the similarity of their embeddings within a shared multimodal space using the CLIP model.
+For Inversion tasks, our evaluation metrics include MSE, LPIPS, SSIM, and PSNR.
+MSE measures the average squared difference between predicted and ground-truth values, quantifying the overall error in pixel intensity.
+LPIPS assesses perceptual similarity between images by comparing deep feature representations extracted from neural networks, aligning with human perception.
+SSIM evaluates image quality by comparing luminance, contrast, and structure to measure the similarity between the reference and reconstructed images.
+PSNR quantifies the ratio between the maximum possible signal value and the power of noise, commonly used to assess image reconstruction quality.
+For video editing, we adopt the VBench Metrics [26].
+The evaluation criteria include Subject Consistency, Motion Smoothness, Aesthetic Quality, and Imaging Quality.
+Subject Consistency measures whether the subject (e.g., a person) remains consistent throughout the video by computing the similarity of DINO features [6] across frames.
+Motion Smoothness assesses the smoothness of motion in the generated video using motion priors from the video frame interpolation model [38].
+Aesthetic Quality evaluates the artistic and visual appeal of each frame as perceived by humans, leveraging the LAION aesthetic predictor [34].
+Imaging Quality examines the level of distortion in the generated frames (e.g., blurring or flickering) based on the MUSIQ image quality predictor [32].
+7 Stylization
+We provide more results on image stylization (Fig.
+9).
+In this task, we perform inversion on the image with a reference style, following the same overall pipeline as in image editing.
+RF-Edit demonstrates superior performance in stylized generation, effectively preserving the style of the source image while generating images that align with the user-provided prompts.
+8 Limitation
+Despite achieving impressive performance across various tasks, our methods occasionally exhibit performance instability, particularly in video editing.
+Additionally, the FLUX and OpenSora models contain a large number of parameters, leading to significant memory consumption, which may limit the applicability of our method on resource-constrained devices.
+To mitigate these issues, we plan to further optimize our method and explore its efficacy on more lightweight models.

### v3

@@
 Extensive experiments across generation, inversion, and editing tasks in both image and video modalities demonstrate the superiority and versatility of our method.
-The source code is available at this URL.
-{strip}[Uncaptioned image]
+Code is available at this URL .
+Machine Learning, ICML
+[…]
+Junfu Pujevinpu@tencent.com
+{strip}
 Figure 1: RF-Solver for downstream tasks in image and video.
-We propose RF-Solver to solve the rectified flow ODE with reduced error, thereby enhancing both sampling quality and inversion-reconstruction accuracy for rectified-flow-based generative models [2, 3].
-Furthermore, we propose RF-Edit, which utilizes the RF-Solver for image and video editing tasks.
+We propose RF-Solver to solve the rectified flow ODE with reduced error, thereby enhancing both sampling quality and inversion-reconstruction accuracy for rectified-flow-based generative models (Black-Forest-Labs, 2024; Zheng et al., 2024).
+Furthermore, we propose RF-Edit, which utilizes the RF-Solver for editing.
 Our methods demonstrate impressive performance across generation, inversion, and editing tasks in both image and video modalities.
 1 Introduction
-Recent advancements of generation methods based on Rectified Flow (RF) [42, 35, 67] have demonstrated exceptional performance in synthesizing high-quality images and videos.
-Different from traditional approaches represented by Stable Diffusion [22, 56], these methods leverage the Diffusion Transformer [51, 78, 73, 63] architecture and implement a straight-line motion system to produce the desired data distribution.
-With these effective designs, FLUX [2] and OpenSora [3] have respectively emerged as one of the state-of-the-art (SOTA) methods in the field of Text-to-Image (T2I) and Text-to-Video (T2V) generation.
-Despite the remarkable success in the fundamental T2I and T2V generation tasks, few studies have explored the performance of RF-based models on various downstream tasks such as inversion-reconstruction [60, 47, 17, 71] and editing [20, 45].
+Recent advancements of generation methods based on Rectified Flow (RF) (Liu et al., 2022a; Lee et al., 2024; Wang et al., 2024b) have demonstrated exceptional performance in synthesizing high-quality images and videos.
+Different from traditional approaches represented by Stable Diffusion (Ho et al., 2020; Rombach et al., 2022), these methods leverage the Diffusion Transformer (Peebles & Xie, 2023; Yang et al., 2024c; Xie et al., 2024; Tang et al., 2024) architecture and implement a straight-line motion system to produce the desired data distribution.
+With these effective designs, FLUX (Black-Forest-Labs, 2024) and OpenSora (Zheng et al., 2024) have respectively emerged as one of the state-of-the-art (SOTA) methods in the field of Text-to-Image (T2I) and Text-to-Video (T2V) generation.
+Despite the remarkable success in the fundamental T2I and T2V generation tasks, few studies have explored the performance of RF-based models on various downstream tasks such as inversion-reconstruction (Song et al., 2021a; Mokady et al., 2023; Guo et al., 2024; Wang et al., 2023c) and editing (Hertz et al., 2022; Meng et al., 2022).
 When directly applying the vanilla RF for inversion, we observe that it fails to faithfully reconstruct the image or video from the source.
-Examples are shown in Fig.
-1 Task 1 and Task 2 (the third row).
-For image inversion, the positions of objects (e.g.
-the cake) and the appearance of individuals (e.g.
-the child) in the reconstructed image significantly diverge from the source image.
+Examples are shown in Figure 1 Task 1 and Task 2 (the third row).
+For image inversion, the positions of objects (e.g., the cake) and the appearance of individuals (e.g., the child) in the reconstructed image significantly diverge from the source image.
 The performance of video inversion is even worse, with noticeable distortions present in the reconstructed video.
-The inaccuracies of inversion and reconstruction would severely constrain the performance of RF models on other inversion-based downstream tasks such as image editing [20, 64, 48, 13, 29] and video editing [41, 33, 15, 59].
+The inaccuracies of inversion and reconstruction would severely constrain the performance of RF models on other inversion-based downstream tasks such as image editing (Hertz et al., 2022; Tumanyan et al., 2023; Nguyen et al., 2024; Duan et al., 2024; Ju et al., 2024) and video editing (Liu et al., 2024; Ku et al., 2024; Fan et al., 2024; Shin et al., 2024).
 In this work, we investigate the aforementioned problem by delving into the inversion and reconstruction process of the RF.
 Specifically, we track the latent at each intermediate timestep during inversion and reconstruction, calculating the Mean Square Error (MSE) between them at corresponding timesteps.
-We observe that significant errors are introduced at each timestep throughout the whole reconstruction process, and their accumulation ultimately results in a considerably deviated output (the red curve in Fig.
-2).
-Based on the definition and inference process of RF [40, 42], we identify that these errors stem from the Ordinary Differential Equation (ODE) solving process [79, 66, 23, 52].
-Specifically, the essence of the inversion and generation process for RF is to derive the solution of RF ODE [42].
+We observe that significant errors are introduced at each timestep throughout the whole reconstruction process, and their accumulation ultimately results in a considerably deviated output (the red curve in Figure 2).
+Based on the definition and inference process of RF (Liu, 2022; Liu et al., 2022a), we identify that these errors stem from the Ordinary Differential Equation (ODE) solving process (Zhang et al., 2024; Wang et al., 2024a; Hong et al., 2024; Peng et al., 2024).
+Specifically, the essence of the inversion and generation process for RF is to derive the solution of RF ODE (Liu et al., 2022a).
 Since this ODE includes terms involving complex neural networks, the solution can only be coarsely approximated by a sampler.
-However, the experiment in Fig.
-2 indicates that the sampler adopted in existing models [2, 3] lacks sufficient precision for the inversion task, causing notable errors to accumulate at each timestep, finally leading to unsatisfactory reconstruction results.
+However, the experiment in Figure 2 indicates that the sampler adopted in existing models (Black-Forest-Labs, 2024; Zheng et al., 2024) lacks sufficient precision for the inversion task, causing notable errors to accumulate at each timestep, finally leading to unsatisfactory reconstruction results.
 Figure 2: Analysis of the inversion-reconstruction process.
@@
 Specifically, we note that the exact formulation of the RF ODE solution can be directly derived using the variation of constants method.
-For the nonlinear component of this solution (i.e., the integral of the neural network), we utilize Taylor expansion for estimation.
+For the nonlinear component of this solution (i.e.,, the integral of the neural network), we utilize Taylor expansion for estimation.
 By employing higher-order Taylor expansion, the ODE can be solved with reduced error, thereby enhancing the performance of RF models.
 RF-Solver is a generic sampler that can be seamlessly integrated into any rectified flow model without additional training or optimization.
-Experimental results demonstrate that RF-Solver not only significantly enhances the accuracy of inversion and reconstruction (the green curve in Fig.
-2), but also improves performance on fundamental tasks such as T2I generation.
+Experimental results demonstrate that RF-Solver not only significantly enhances the accuracy of inversion and reconstruction (the green curve in Figure 2), but also improves performance on fundamental tasks such as T2I generation.
 Building upon this, we propose RF-Edit to leverage RF-Solver in editing tasks.
 Real-world image and video editing require the model to make precise modifications to a source image/video while maintaining its overall structure unchanged, presenting greater challenges than reconstruction.
-In this scenario, it is inadequate to solely rely on the inverted noises as prior knowledge for editing, which could lead to edited results being excessively influenced by the target prompt, diverging significantly from the original source [20, 64].
+In this scenario, it is inadequate to solely rely on the inverted noises as prior knowledge for editing, which could lead to edited results being excessively influenced by the target prompt, diverging significantly from the original source (Hertz et al., 2022; Tumanyan et al., 2023).
 Addressing this problem, RF-Edit stores the $\mathcal{V}$ (value) feature in the self-attention layers at several timesteps during inversion.
 These features are used to replace the corresponding features in the denoising process.
-Practically, we design two specific sub-modules for RF-Edit, respectively leveraging the DiT structure of FLUX [2] and OpenSora [3] as the backbones for image and video editing.
+Practically, we design two specific sub-modules for RF-Edit, respectively leveraging the DiT structure of FLUX (Black-Forest-Labs, 2024) and OpenSora (Zheng et al., 2024) as the backbones for image and video editing.
 With the effective design of RF-Edit, it demonstrates superior performance in both image and video domains, outperforming various SOTA methods.
-Our core contributions are summarized as follows:
-  • We propose RF-Solver, a training-free sampler that significantly reduces errors in the inversion and reconstruction processes of rectified-flow models.
-  • We present RF-Edit, which leverages RF-Solver for image and video editing. RF-Edit effectively preserves the structural integrity of the source image/video while delivering high-quality results.
-  • Extensive experiments across a range of tasks demonstrate the efficacy and versatility of our method.
 2 Related Work
 2.1 Inversion
-Inversion maps the real visual data, i.e.
-image and video, to representations in noise space, which is the reverse process of generation.
-The representative method, DDIM inversion [60, 61], adds predicted noise recursively at each forward step.
-Many efforts [14, 65, 47, 57, 46, 27, 46, 44, 19] have been made to mitigate the discretization error in DDIM inversion.
+Inversion maps the real visual data, i.e., image and video, to representations in noise space, which is the reverse process of generation.
+The representative method, DDIM inversion (Song et al., 2021a, b), adds predicted noise recursively at each forward step.
+Many efforts (Elarabawy et al., 2022; Wallace et al., 2023; Mokady et al., 2023; Rout et al., 2024a; Miyake et al., 2023; Lu et al., 2022) have been made to mitigate the discretization error in DDIM inversion.
 Despite the effectiveness of inversion in diffusion models, the exploration of inversion in SOTA rectified flow models like FLUX and OpenSora is limited.
-RF-prior [76] uses the score distillation to invert the image while it requires many optimizing steps.
-More recently, [58] introduces an additional vector field conditioned on the source image to improve the inversion.
+RF-prior (Yang et al., 2024b) uses the score distillation to invert the image while it requires many optimizing steps.
+More recently, (Rout et al., 2024b) introduces an additional vector field conditioned on the source image to improve the inversion.
 However, the error from the original vector field of rectified flow still persists, which would limit the performance of such method on various downstream tasks.
@@
 2.2 Image and Video Editing
-Training-free methods for image and video editing [25, 62] have gained increasing popularity for their efficiency and effectiveness.
-Existing image editing methods focus on prompt refinement [55, 69], attention-sharing mechanism [20, 50, 18, 64], mask guidance [4, 10, 24, 37, 19], and noise initialization [5, 77].
+Training-free methods for image and video editing (Huang et al., 2024a; Sun et al., 2024) have gained increasing popularity for their efficiency and effectiveness.
+Existing image editing methods focus on prompt refinement (Ravi et al., 2023; Wang et al., 2023a), attention-sharing mechanism (Hertz et al., 2022; Parmar et al., 2023; Cao et al., 2023; Tumanyan et al., 2023), mask guidance (Avrahami et al., 2023; Couairon et al., 2022; Huang et al., 2023), and noise initialization (Brack et al., 2023; Yang et al., 2023b).
 Video editing introduces additional complexities in maintaining temporal consistency, making it a more challenging task.
-Existing video editing methods focus on attention injection [53, 70, 41], motion guidance [9, 16, 75, 68], latent manipulation [81, 74, 30, 8], and canonical representation [7, 36, 49, 31].
+Existing video editing methods focus on attention injection (Qi et al., 2023; Wang et al., 2023b; Liu et al., 2024), motion guidance (Cong et al., 2023; Geyer et al., 2023; Yang et al., 2024a; Wang et al., 2024c), latent manipulation (Zhang et al., 2023; Yang et al., 2023a; Kara et al., 2024; Chen et al., 2023), and canonical representation (Chai et al., 2023; Lee et al., 2023; Ouyang et al., 2024; Kasten et al., 2021).
 To date, the editing performance of RF-based diffusion transformers has remained largely under-explored.
-Although [58] employs FLUX [2] for image editing, its performance is limited to simple tasks such as stylization and face editing while often failing to effectively maintain the structural information of source images.
+Although (Rout et al., 2024b) employs FLUX (Black-Forest-Labs, 2024) for image editing, its performance is limited to simple tasks such as stylization and face editing while often failing to effectively maintain the structural information of source images.
 Moreover, currently there is no research exploring the video editing capabilities of RF-based models.
@@
 3.1 Preliminaries
-Rectified Flow (RF) [43] facilitates the transition between the real data distribution $\pi_{0}$ and Gaussian Noises distribution $\pi_{1}$ along a straight path.
+Rectified Flow (RF) (Liu et al., 2022b) facilitates the transition between the real data distribution $\pi_{0}$ and Gaussian Noises distribution $\pi_{1}$ along a straight path.
 This is achieved by learning a forward-simulating system defined by the ODE: $d\boldsymbol{Z}_{t}=v(\boldsymbol{Z}_{t},t)dt,t\in[0,1]$, which maps $\boldsymbol{Z}_{1}\in\pi_{1}$ to $\boldsymbol{Z}_{0}\in\pi_{0}$.
 In practice, the velocity field $v$ is parameterized by a neural network $\boldsymbol{v}_{\theta}$.
-During training, given empirical observations of two distributions $\boldsymbol{X}_{0}\sim\pi_{0}$, $\boldsymbol{X}_{1}\sim\pi_{1}$ and $t\in[0,1]$, the forward process (i.e.
-adding noise) of rectified flow is defined by a simple linear combination: $\boldsymbol{X}_{t}=t\boldsymbol{X}_{1}+(1-t)\boldsymbol{X}_{0}$.
+During training, given empirical observations of two distributions $\boldsymbol{X}_{0}\sim\pi_{0}$, $\boldsymbol{X}_{1}\sim\pi_{1}$ and $t\in[0,1]$, the forward process (i.e., adding noise) of rectified flow is defined by a simple linear combination: $\boldsymbol{X}_{t}=t\boldsymbol{X}_{1}+(1-t)\boldsymbol{X}_{0}$.
 The differential form of the equation is given by: $d\boldsymbol{X}_{t}=(\boldsymbol{X}_{1}-\boldsymbol{X}_{0})dt$.
@@
 $\boldsymbol{Z}_{t_{i-1}}=\boldsymbol{Z}_{t_{i}}+(t_{i-1}-t_{i})\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i}).$ (2)
-The RF model can generate high-quality images in much fewer timesteps compared to DDPM [22], owing to the nearly linear transition trajectory established during training.
-With these effective designs, RF model illustrates great potential in the field of T2I and T2V generation [2, 3].
+The RF model can generate high-quality images in much fewer timesteps compared to DDPM (Ho et al., 2020), owing to the nearly linear transition trajectory established during training.
+With these effective designs, RF model illustrates great potential in the field of T2I and T2V generation (Black-Forest-Labs, 2024; Zheng et al., 2024).
 3.2 RF-Solver
@@
 However, when applied to inversion and reconstruction tasks, we observe significant error accumulation at each timestep.
-This results in reconstructions that diverge notably from the original image (see Fig.
-2).
-This severely limits the performance of RF models in various inversion-based downstream tasks [20, 66].
-Delving into this problem, we identify that the errors stem from the process of estimating the approximate solution for the rectified flow ODE [67, 40], which is formulated by Eq.
-2 in existing methods [2, 3].
-Consequently, obtaining more precise solutions for the ODE would effectively mitigate these errors, leading to improved reconstruction quality.
+This results in reconstructions that diverge notably from the original image (see Figure 2).
+This severely limits the performance of RF models in various inversion-based downstream tasks (Hertz et al., 2022; Wang et al., 2024a).
+Delving into this problem, we identify that the errors stem from the process of estimating the approximate solution for the rectified flow ODE (Wang et al., 2024b; Liu, 2022), which is formulated by Equation 2 in existing methods (Black-Forest-Labs, 2024; Zheng et al., 2024).
+Consequently, obtaining more precise solutions for the ODE would effectively mitigate these errors, leading to improved performance.
 Based on this analysis, we start by carefully examining the differential form of the Rectified flow: ${d\boldsymbol{Z}_{t}}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t},t)dt$.
@@
 where $\boldsymbol{v}^{(k)}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i})=\frac{\text{d}^{k}\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i})}{\text{d}t^{k}}$, denoting the $k$-order derivative of $\boldsymbol{v}_{\theta}$ and $\mathcal{O}$ denotes higher-order infinitesimals.
-Substituting Eq.
-4 into the integral term yields:
+Substituting Equation 4 into the integral term yields:
 […]
@@
 […]
-Substituting Eq. 6 and Sec. 3.2 into Eq. 3, we derive the $n$-th order solution of Rectified flow ODE:
+Substituting Equation 6 and Section 3.2 into Equation 3, we derive the $n$-th order solution of Rectified flow ODE:
 $\boldsymbol{Z}_{t_{i-1}}=\boldsymbol{Z}_{t_{i}}$ $+\sum\limits_{k=0}^{n-1}\frac{({t_{i-1}}-{t_{i}})^{k+1}}{(k+1)!}{\boldsymbol{v}}^{(k)}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$
@@
 where $h_{i}:=t_{i-1}-t_{i}$.
-Sec.
-3.2 indicates that to estimate $\boldsymbol{Z}_{t_{i-1}}$, we need to obtain the $k$-th order derivatives $\{\boldsymbol{v}^{(k)}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})\}$ for $k\in\{0,\cdots,n-1\}$.
-When $n=1$, the formula reduces to the standard rectified flow (i.e., Eq. 2). In our experiments, we find that setting $n=2$ effectively mitigates the errors, yielding:
+Section 3.2 indicates that to estimate $\boldsymbol{Z}_{t_{i-1}}$, we need to obtain the $k$-th order derivatives $\{\boldsymbol{v}^{(k)}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})\}$ for $k\in\{0,\cdots,n-1\}$.
+Considering $n=1$, the formula reduces to the standard rectified flow (i.e.,, Equation 2). In our experiments, we find that setting $n=2$ effectively mitigates the errors, yielding:
 $\boldsymbol{Z}_{t_{i-1}}=\boldsymbol{Z}_{t_{i}}$ $+({t_{i-1}}-{t_{i}})\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i})$
@@
 We design two sub-modules for applying RF-Edit to (a).
-Image editing with FLUX [2] and (b).
-Video editing with OpenSora [3].
+Image editing with FLUX (Black-Forest-Labs, 2024) and (b).
+Video editing with OpenSora (Zheng et al., 2024).
 Note that for FLUX, there are multiple Double Blocks, followed by multiple Single Blocks.
@@
 For simplicity, only one block of each type is depicted in the figure.
-Note that $\boldsymbol{v}_{\theta}^{(1)}$ in Sec.
-3.2 is the first-order derivative of the network prediction term $\boldsymbol{v}_{\theta}$, which cannot be analytically derived due to the complex architecture of the neural network.
+Note that $\boldsymbol{v}_{\theta}^{(1)}$ in Section 3.2 is the first-order derivative of the network prediction term $\boldsymbol{v}_{\theta}$, which cannot be analytically derived due to the complex architecture of the neural network.
 To estimate this term, we first obtain the network prediction $\hat{\boldsymbol{v}}_{t_{i}}$ at the timestep $t_{i}$, i.e., $\hat{\boldsymbol{v}}_{t_{i}}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$.
-Then we step forward a small timestep $\Delta t=\frac{1}{2}(t_{i-1}-t_{i})$, and update the latents to obtain $\boldsymbol{Z}_{{t_{i}}+\Delta t}=\boldsymbol{Z}_{t_{i}}+\Delta t\cdot\hat{\boldsymbol{v}}_{t_{i}}$.
+Then we step forward a small timestep $\Delta t$ (which is set to $0.01$ in experiments), and update the latents to obtain $\boldsymbol{Z}_{{t_{i}}+\Delta t}=\boldsymbol{Z}_{t_{i}}+\Delta t\cdot\hat{\boldsymbol{v}}_{t_{i}}$.
 Subsequently, we calculate an additional prediction of the network at the timestep $t_{i}+\Delta t$, i.e., $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}+\Delta t},{t_{i}+\Delta t})$.
 With $\hat{\boldsymbol{v}}_{t_{i}}$ and $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}$, the first-order derivative of ${\boldsymbol{v}}_{\theta}$ at the timestep $t_{i}$ can be estimated as: $\boldsymbol{v}_{\theta}^{(1)}(\boldsymbol{Z}_{t_{i}},t_{i})=\frac{\hat{\boldsymbol{v}}_{t_{i}+\Delta t}-\hat{\boldsymbol{v}}_{t_{i}}}{\Delta t}$.
-Substituting this formulation into Sec.
-3.2 results in the practical implementation of the RF-Solver algorithm.
+Substituting this formulation into Section 3.2 results in the practical implementation of the RF-Solver algorithm.
 The complete sampling process for RF-Solver is presented in Algorithm 1.
-Algorithm 1 Sampling process of RF-Solver
-Input:
-  $\boldsymbol{v}_{\theta}$triangleright Velocity function
-  $t=[t_{N},\ldots,t_{0}]$triangleright Time steps
-  $\boldsymbol{Z}_{t_{N}}\sim\mathcal{N}(0,I)$triangleright Initial Gaussian Noise
-For $i=N$ to $1$ do
-  $\Delta t_{i}\leftarrow\frac{1}{2}(t_{i-1}-t_{i})$
-  $\boldsymbol{\hat{v}}_{t_{i}}\leftarrow\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i})$
-  $\boldsymbol{Z}_{t_{i}+\Delta t_{i}}\leftarrow\boldsymbol{Z}_{t_{i}}+\Delta t_{i}\boldsymbol{\hat{v}}_{t_{i}}$
-  $\boldsymbol{\hat{v}}_{t_{i}+\Delta t_{i}}\leftarrow\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}+\Delta t_{i}},t_{i}+\Delta t_{i})$
-  $\boldsymbol{v}_{t_{i}}^{(1)}\leftarrow({\boldsymbol{\hat{v}}_{t_{i}+\Delta t_{i}}-\boldsymbol{\hat{v}}_{t_{i}}})/{\Delta t_{i}}$triangleright Calculating the Derivatives
-  $\boldsymbol{Z}_{t_{i-1}}\leftarrow\boldsymbol{Z}_{t_{i}}+(t_{i-1}-t_{i})\boldsymbol{\hat{v}}_{t_{i}}+\frac{1}{2}(t_{i-1}-t_{i})^{2}\boldsymbol{v}_{t_{i}}^{(1)}$
-Output: $\boldsymbol{Z}_{0}$
 Obtaining the sampling form of RF-Solver, we further derive its inversion form.
 Inversion maps data back into noise, which reverses the sampling process.
-Following previous methods for DDIM inversion [60, 12], the ODE process can be directly reversed in the limit of small steps.
-Based on this assumption, the inversion process of RF-Solver (Sec.
-3.2) can be directly derived as:
+Following previous methods for DDIM inversion (Song et al., 2021a; Dhariwal & Nichol, 2021), the ODE process can be directly reversed in the limit of small steps.
+Based on this assumption, the inversion process of RF-Solver can be derived as:
 $\widetilde{\boldsymbol{Z}}_{t_{i+1}}=\widetilde{\boldsymbol{Z}}_{t_{i}}$ $+({t_{i+1}}-{t_{i}})\boldsymbol{v}_{\theta}(\widetilde{\boldsymbol{Z}}_{t_{i}},t_{i})$
@@
 where $\widetilde{\boldsymbol{Z}}_{t_{i}}$ and $\widetilde{\boldsymbol{Z}}_{t_{i+1}}$ denotes the latents during inversion.
-Through the high order expansion, the error of the ODE solution in each timestep is reduced from $\mathcal{O}\big{(}(h_{i})^{2}\big{)}$ to $\mathcal{O}\big{(}(h_{i})^{3}\big{)}$, leading to improved performance, particularly in inversion and reconstruction (see Fig.
-2).
-Beyond inversion and reconstruction, RF-Solver can also be applied to any RF-based model (such as FLUX [2] and OpenSora [3]) for other tasks such as sampling and editing, enhancing performance without requiring additional training.
+Through the high order expansion, the error of the ODE solution in each timestep is reduced from $\mathcal{O}\big{(}(h_{i})^{2}\big{)}$ to $\mathcal{O}\big{(}(h_{i})^{3}\big{)}$, significantly facilitating inversion and reconstruction process (see Figure 2).
+Beyond this, RF-Solver can also be applied to any RF-based model (such as FLUX (Black-Forest-Labs, 2024) and OpenSora (Zheng et al., 2024)) for other tasks such as sampling and editing, enhancing performance without requiring additional training.
 3.3 RF-Edit
@@
 However, directly applying RF-Solver during the inversion and denoising stages may cause the model to be overly influenced by the target prompt, resulting in unintended modifications in other regions of the source image or video.
-Similar issues are common across various existing editing methods [58, 20, 64].
+Similar issues are common across various existing editing methods (Rout et al., 2024b; Hertz et al., 2022; Tumanyan et al., 2023).
 To address this problem, we propose RF-Edit, which builds upon the diffusion transformer architecture.
@@
 The proposed RF-Edit framework enables high-quality editing while effectively preserving the structural information of the source image/video.
-Building on this concept, we design two sub-modules for RF-Edit, specifically tailored for image editing and video editing (Fig.
-3).
-For image editing, RF-Edit employs FLUX [2] as the backbone, which comprises several double blocks and single blocks.
+Building on this concept, we design two sub-modules for RF-Edit, specifically tailored for image editing and video editing (Figure 3).
+For image editing, RF-Edit employs FLUX (Black-Forest-Labs, 2024) as the backbone, which comprises several double blocks and single blocks.
 Double blocks independently modulate text and image features, while single blocks concatenate these features for unified modulation.
 In this architecture, RF-Edit shares features within the single blocks, as they capture information from both the source image and the source prompt, enhancing the ability of the model to preserve the structural information of the source image.
-For video editing, RF-Edit employs OpenSora [3] as the backbone.
+For video editing, RF-Edit mainly employs OpenSora (Zheng et al., 2024) as the backbone.
 The DiT blocks in OpenSora include spatial attention, temporal attention, and text cross-attention.
@@
 4.1 Setup
-Baselines.
-We use the vanilla Rectified Flow sampler (Euler Sampler) as the primary baseline for all tasks.
-Besides, for sampling and inversion, we compare our method with DPM-Solver++ [44] and the Heun sampler (a 2-order ODE solver).
-For image editing, we compare our method with P2P [20], DiffEdit [11], SDEdit [45], PnP [64], Pix2pix [50] and RF-Inversion [58].
-For video editing tasks, we compare our method with FateZero [53], FLATTEN [9], COVE [68], RAVE [30], Tokenflow [16].
-Detailed experimental settings of baselines are provided in the Appendix.
-Figure 5: Qualitative results of text-to-image generation. By employing the RF-Solver, the model is able to generate images with higher quality (the first row) than baselines (the second row).
-Implementation Details.
-In the experiment, we adopt the guidance-distilled variant of FLUX [2] for image tasks and OpenSora [3] for video tasks.
+We implement our method respectively on FLUX (Black-Forest-Labs, 2024) and OpenSora (Zheng et al., 2024).
+In the experiment, we adopt the guidance-distilled variant of FLUX (Black-Forest-Labs, 2024) for image tasks and OpenSora (Zheng et al., 2024) for video tasks.
 The derivative computation in RF-Solver requires an additional forward pass, resulting in the network needing to forward twice at each timestep.
 As a result, when comparing our method with the Rectified Flow baselines, we set the number of timesteps for the vanilla Rectified Flow to be twice that of our method to ensure a fair comparison under the same number of function evaluations (NFE).
-More information is provided in the Appendix.
-Evaluation Metrics For text-to-image sampling, we randomly select 10k images from the MSCOCO validation set [39] and report the FID [21] and CLIP Score [54].
-For the inversion and reconstruction task, we report the Mean Square Error (MSE), LPIPS [80], SSIM [72], and PSNR [28].
-For image editing tasks, we report the CLIP Score [54] and LPIPS [80].
-For video editing tasks, we adopt the metric proposed by [26], including Subject Consistency (SC), Motion Smoothness (MS), Aesthetic Quality (AQ), and Imaging Quality (IQ).
-Detailed explanations of these metrics are provided in the Appendix.
+More detailed information for experiment setup is provided in Appendix B.
+Table 1: Quantitative results on text-to-image generation. RF-Solver outperforms several baselines.
+DPMSolver++ RF RF-Heun Ours
+FID ($\downarrow$) 24.63 25.33 24.40 23.98
+CLIP Score ($\uparrow$) 30.62 31.01 31.03 31.09
 4.2 Text-to-image Sampling
-We compare the performance of our method with DPMSolver++ [44], the vanilla RF sampler, and Heun sampler on the text-to-image generation task.
-Both the quantitative (Tab.
-1) and qualitative results (Fig.
-5) demonstrate the superior performance of RF-Solver in fundamental T2I generation tasks, producing higher-quality images that align more closely with human cognition.
+We compare the performance of our method with DPMSolver++ (Lu et al., 2022), the vanilla RF sampler, and Heun sampler on the text-to-image generation task.
+Both the quantitative (Table 1) and qualitative results (Figure 10) demonstrate the superior performance of RF-Solver in fundamental T2I generation tasks, producing higher-quality images that align more closely with human cognition.
 Figure 5: Qualitative comparison of image editing.
 With RF-Solver and feature-sharing mechanism in RF-Edit, our method can successfully handle various kinds of image editing cases, outperforming the previous SOTA methods.
-Zoom in for the best views.
-DPMSolver++ RF RF-Heun Ours
-FID ($\downarrow$) 24.63 25.33 24.40 24.03
-CLIP Score ($\uparrow$) 30.62 31.01 31.03 31.09
-Table 1: Quantitative results on text-to-image generation. RF-Solver outperforms several baselines.
 4.3 Inversion and Reconstruction
@@
 Quantitative Comparison.
-The quantitative comparisons (Tab.
-2) are conducted to reflect the similarity between the source and reconstruction results.
+The quantitative comparisons (Table 2) are conducted to reflect the similarity between the source and reconstruction results.
 Our method demonstrates superior performance across all four metrics compared with the vanilla RF sampler and Heun sampler.
-Mehtod MSE ($\downarrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$)
+Table 2: Quantitative results on inversion and reconstruction. Our method significantly improves the accuracy of reconstruction for both images and videos.
+Method MSE ($\downarrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$)
 image RF 0.0268 0.6253 0.7626 28.28
 RF-Heun 0.0117 0.4696 0.8924 29.67
-Ours 0.0094 0.4242 0.9271 29.83
+Ours 0.0092 0.4239 0.9276 29.89
 video RF 0.0206 0.4159 0.8134 18.12
 RF-Heun 0.0156 0.3554 0.8711 18.29
-Ours 0.0139 0.3299 0.8805 18.32
-Table 2: Quantitative results on inversion and reconstruction. Our method significantly improves the accuracy of reconstruction for both images and videos.
+Ours 0.0134 0.3287 0.8812 18.41
 Qualitative Comparison.
 RF-Solver effectively reduces the error in the solution of RF ODE, thereby increasing the accuracy of the reconstruction.
-As illustrated in Fig.
-4(a), the image reconstruction results using vanilla rectified flow exhibit noticeable drift from the source image, with significant alterations to the appearance of subjects in the image.
-For video reconstruction, as shown in Fig.
-4(b), the baseline reconstruction results suffer from distortion.
+As illustrated in Figure 4(a), the image reconstruction results using vanilla rectified flow exhibit noticeable drift from the source image, with significant alterations to the appearance of subjects in the image.
+For video reconstruction, as shown in Figure 4(b), the baseline reconstruction results suffer from distortion.
 In contrast, RF-Solver significantly alleviates these issues, achieving more satisfactory results.
-Figure 7: Qualitative comparison of video editing.
-The first video comprises 200 frames with a resolution of $512\times 512$, while the second video contains 60 frames with a resolution of $1024\times 768$ (frames are compressed for a neat layout in the figure).
 4.4 Editing
 We conduct experiments to evaluate the image and video editing performance of our method.
-Image editing usually involves replacing the subject in the image with another one, adding new items, and global editing.
-For the first two types of editing, the background of the source image is expected to remain unchanged after editing.
-For global editing such as style transfer, the overall structure of the source image is expected to remain unchanged.
-Recent mainstream video editing methods usually focus on replacing the subjects and performing global editing for the source video.
-Quantitative Comparison.
-In image editing, Our method outperforms all other methods in CLIP score (Tab.
-3), indicating that the edited images align well with the user-provided prompts.
-For LPIPS, it is noted that PnP [64] has a much lower value than all other methods.
-Based on the qualitative results (Fig.
-6), it can be seen that PnP is only suitable for editing cases that do not significantly modify the structure or shape of the source image (such as changing red roses into yellow sunflowers).
-It fails in the case of shape editing, resulting in an image very similar to the source.
-Consequently, although PnP has the lowest LPIPS score, its CLIP score is the lowest.
+For image editing, we compare our method with P2P (Hertz et al., 2022), DiffEdit (Couairon et al., 2023), SDEdit (Meng et al., 2022), PnP (Tumanyan et al., 2023), Pix2pix (Parmar et al., 2023) and RF-Inversion (Rout et al., 2024b).
+For video editing tasks, we compare our method with FateZero (Qi et al., 2023), FLATTEN (Cong et al., 2023), COVE (Wang et al., 2024c), RAVE (Kara et al., 2024), Tokenflow (Geyer et al., 2023).
+Table 3: Quantitative results of image editing. RF-Edit effectively edits the images according to the prompts while preserving the integrity of unrelated regions.
 P2P DiffEdit SDEdit PnP Pix2Pix RF-Inv Ours
@@
 CLIP Score ($\uparrow$) 30.70 32.68 31.61 30.58 32.33 33.02 33.66
-Table 3: Quantitative results of image editing. RF-Edit effectively edit the images according to the prompts while preserving the integrity of unrelated regions.
-For video editing, RF-Edit achieves higher scores on the popular VBench [26] metrics (Tab.
-4).
-The results illustrate that our method successfully maintains temporal consistency while demonstrating superior visual quality.
+Quantitative Comparison.
+In image editing, Our method outperforms all other methods in CLIP score (Table 3), indicating that the edited images align well with the user-provided prompts.
+For LPIPS, it is noted that PnP (Tumanyan et al., 2023) has a much lower value than all other methods.
+Based on the qualitative results (Figure 5), it can be seen that PnP is only suitable for editing cases that do not significantly modify the structure or shape of the source image (such as changing red roses into yellow sunflowers).
+It fails in the case of shape editing, resulting in an image very similar to the source.
+Consequently, although PnP has the lowest LPIPS score, its CLIP score is the lowest.
+For video editing, RF-Edit achieves higher scores on VBench (Huang et al., 2024b) metrics (Table 4).
+The results illustrate that our method successfully maintains temporal consistency while demonstrating superior quality.
+Figure 6: Qualitative comparison of video editing.
+The first video comprises 200 frames with a resolution of $512\times 512$, while the second video contains 60 frames with a resolution of $1024\times 768$ (frames are compressed for a neat layout).
+Table 4: Quantitative results of video editing. RF-Edit outperforms several previous SOTA video editing methods.
 FateZero Flatten COVE RAVE Tokenflow Ours
@@
 IQ ($\uparrow$) 0.6898 0.7024 0.7163 0.6917 0.7128 0.7207
-Table 4: Quantitative results of video editing. RF-Edit outperforms several previous SOTA video editing methods.
 Qualitative Comparison.
-For image editing, we compare the performance of our method with several baselines across different types of editing tasks (Fig.
-6).
+For image editing, we compare the performance of our method with several baselines across different types of editing requirements including adding, replacing, and stylization (Figure 5 and Figure 8).
 The baseline methods often suffer from background changes or fail to perform the desired edits.
 In contrast, our method demonstrates satisfying performance, effectively achieving a balanced trade-off between the fidelity to the target prompt and the preservation of the source image.
-For video editing, we primarily evaluate the performance of our method on long videos (200 frames) and high-resolution videos ($1280\times 768$).
-The qualitative results are shown in Fig.
-7.
+The qualitative results for video editing are shown in Figure 6.
 RF-Edit illustrates impressive performance in handling complicated editing cases (e.g., modifying the leftmost lion among three lions into a white polar bear and changing the other two small lions into orange tiger cubs), whereas all other baseline methods fail in this scenario.
-RF-Edit also demonstrates strong performance in global editing tasks, such as transforming scenes into autumn.
-Metric RF RF-Solver-2 RF-Solver-3
-Sampling FID ($\downarrow$) 25.33 24.03 23.96
-CLIP Score ($\uparrow$) 31.01 31.09 31.09
-Inversion MSE ($\downarrow$) 0.0268 0.0094 0.0131
-LPIPS ($\downarrow$) 0.6253 0.4242 0.4817
-Editing LPIPS ($\downarrow$) 0.1524 0.1494 0.1503
-CLIP Score ($\uparrow$) 32.97 33.66 33.18
-Table 5: Ablation study on the Taylor Expansion order. We select the 2-order expansion (i.e. RF-Solver-2) for various downstream tasks due to its effectiveness and simplicity.
+Besides, HunyuanVideo (Kong et al., 2024) has recently demonstrated strong performance in text-to-video generation.
+Thanks to the generality of our method, it can be implemented on HunyuanVideo for video editing.
+More qualitative results are shown in Figure 9.
 4.5 Ablation Study
 We conduct ablation studies to illustrate the effectiveness of RF-Solver and RF-Edit.
-Without loss of generality, these ablation studies are performed on the image tasks using FLUX [2] as the base model.
+Without loss of generality, these ablation studies are performed on the image tasks using FLUX (Black-Forest-Labs, 2024) as the base model.
 Taylor Expansion Order of RF-Solver.
-We investigated the impact of the Taylor expansion order in RF-Solver (Tab.
-5) under the same NFE across different orders.
-The second-order expansion demonstrated a significant improvement across various tasks compared to the first-order expansion (i.e., the vanilla rectified flow).
+We investigated the impact of the Taylor expansion order in RF-Solver (Table 5) under the same NFE across different orders.
+Compared to the first-order expansion (i.e., the vanilla rectified flow), the second-order expansion demonstrates a significant improvement across various tasks.
 However, higher-order expansions do not yield further enhancements.
@@
 Moreover, computing the higher-order derivatives of $\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$ substantially increases the complexity of the algorithm, posing challenges for practical applications.
-Consequently, we predominantly employed second-order expansion in our experiments.
-Figure 8: Ablation study of feature-sharing step in RF-Edit.
-A too-small feature-sharing step results in the inconsistency between source and target images.
-Conversely, a too-large feature-sharing step can lead to the failure of editing.
+Consequently, we employ second-order expansion (i.e., RF-Solver-2 in Table 5) for various downstream tasks due to its effectiveness and simplicity.
+Table 5: Ablation study on the Taylor Expansion order.
+Here, RF implies the vanilla Rectified Flow without the proposed RF-Solver algorithm.
+All of the experiments for editing in the table apply the proposed feature-sharing mechanism for better results.
+Metric RF RF-Solver-2 RF-Solver-3
+Sampling FID ($\downarrow$) 25.33 23.98 23.94
+CLIP Score ($\uparrow$) 31.01 31.09 31.09
+Inversion MSE ($\downarrow$) 0.0268 0.0092 0.0131
+LPIPS ($\downarrow$) 0.6253 0.4239 0.4817
+Editing LPIPS ($\downarrow$) 0.1524 0.1494 0.1503
+CLIP Score ($\uparrow$) 32.97 33.66 33.18
 Feature Sharing Steps of RF-Edit.
 RF-Edit leverages feature sharing to maintain the structural consistency between original images and edited images.
-However, an excessive number of feature-sharing steps may result in the edited output being overly similar to the source image, ultimately undermining the intended editing objectives (Fig.
-8).
+However, an excessive number of feature-sharing steps may result in the edited output being overly similar to the source image, ultimately undermining the intended editing objectives (Figure 7).
 To investigate the impact of feature-sharing steps on editing results, we incrementally increase the number of feature-sharing steps applied to the same image.
@@
 Additionally, we can customize the sharing step for each image to identify the most satisfying outcome.
+Figure 7: Ablation study of feature-sharing step in RF-Edit.
+The second row is the results produced solely by RF-Solver, without the proposed feature-sharing mechanism.
+Feature-sharing can significantly enhance the consistency between source and target images, while a too-large feature-sharing step may lead to the failure of editing.
 5 Conclusion
@@
 Extensive experiments demonstrate the versatility and effectiveness of our method.
-Supplementary Material
-6 Experimental Settings
-6.1 Baselines and Implementation Details
+Acknowledgments
+This work was partly supported by Shenzhen Key Labora-tory of next generation interactive media innovative tech-nology (No:ZDSYS20210623092001004) and National Key Laboratory of Human-Machine Hybrid Augmented Intelligence, Xi’an Jiaotong University (No.
+HMHAI202410).
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix
+Appendix A Pesudo Code of RF-Solver Algorithm.
+Algorithm 1 Sampling process of RF-Solver
+Input:
+  $\boldsymbol{v}_{\theta}$triangleright Velocity function
+  $t=[t_{N},\ldots,t_{0}]$triangleright Time steps
+  $\boldsymbol{Z}_{t_{N}}\sim\mathcal{N}(0,I)$triangleright Initial Gaussian Noise
+For $i=N$ to $1$ do
+  $\boldsymbol{\hat{v}}_{t_{i}}\leftarrow\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},t_{i})$
+  $\boldsymbol{Z}_{t_{i}+\Delta t_{i}}\leftarrow\boldsymbol{Z}_{t_{i}}+\Delta t_{i}\boldsymbol{\hat{v}}_{t_{i}}$
+  $\boldsymbol{\hat{v}}_{t_{i}+\Delta t_{i}}\leftarrow\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}+\Delta t_{i}},t_{i}+\Delta t_{i})$
+  $\boldsymbol{v}_{t_{i}}^{(1)}\leftarrow({\boldsymbol{\hat{v}}_{t_{i}+\Delta t_{i}}-\boldsymbol{\hat{v}}_{t_{i}}})/{\Delta t_{i}}$triangleright Calculating the Derivatives
+  $\boldsymbol{Z}_{t_{i-1}}\leftarrow\boldsymbol{Z}_{t_{i}}+(t_{i-1}-t_{i})\boldsymbol{\hat{v}}_{t_{i}}+\frac{1}{2}(t_{i-1}-t_{i})^{2}\boldsymbol{v}_{t_{i}}^{(1)}$
+Output: $\boldsymbol{Z}_{0}$
+Appendix B Experimental Settings
+B.1 Baselines and Implementation Details
 Text-to-Image Generation.
@@
 We compare the performance of our methods among RF with the vanilla sampler and the Heun sampler.
-For image inversion, we also use the images from the MS-COCO validation set.
-For video inversion, we select videos from social media platforms such as TikTok and other publicly available sources.
-We have observed the quality of the text prompts significantly influence the quality of inversion.
+For image inversion, similar to Text-to-Image generation, we use images and captions from the MS-COCO validation set.
+For video inversion, we select about 40 videos from social media platforms such as TikTok and other publicly available sources.
+We have observed the quality of the text prompts significantly influences the quality of inversion.
 Consequently, we employ GPT-4o to generate detailed captions for both images and videos, which are then used in the inversion tasks.
 The total NFE for generating one image/video is set to 50 for both our method and baselines.
+Figure 8: Stylized Generation.
+Figure 9: More video editing results.
 Editing.
-For image editing, we compare our methods with RF-inversion and several diffusion-based editing methods.
-For RF-inversion, we adopt the implementation in ComfyUI [1].
-For other baselines, we use their implementation from diffusers.
-For each baseline, we adjust the relevant hyper-parameters to achieve optimal results.
-For video editing, we use the official codes of all the baseline methods and tune the hyper-parameters to achieve satisfactory results.
 For image editing, we share the features of the last 19 single blocks in FLUX.
@@
 We adjust the hyper-parameter of feature-sharing steps to achieve better results for both image and video editing.
-Figure 9: Stylization Results.
-6.2 Evaluation Metrics
+For image editing, we use over 300 images for quantitative comparison, which both include real images (obtained from public social media and the DIV2K dataset) and generated images.
+For each image, we use GPT-4o to generate the source prompt and manually refine the generated prompt.
+There are 2 3̃ target prompts for different requirements of editing including adding, replacing, and stylization for each source image.
+We compare our methods with RF-inversion and several diffusion-based editing methods.
+For RF-inversion, we adopt the implementation in ComfyUI (com, ).
+For other baselines, we use their implementation from diffuser and adjust the relevant hyper-parameters to achieve optimal results.
+For video editing, the data preparation mainly follows previous works COVE (Wang et al., 2024c).
+We use the official codes of all the baseline methods and tune the hyper-parameters to achieve satisfactory results.
+Figure 10: Qualitative results of text-to-image generation. By employing the RF-Solver, the model is able to generate images with higher quality than baselines.
+B.2 Evaluation Metrics
 For text-to-image sampling, we report Fréchet Inception Distance (FID) and CLIP Scores.
@@
 PSNR quantifies the ratio between the maximum possible signal value and the power of noise, commonly used to assess image reconstruction quality.
-For video editing, we adopt the VBench Metrics [26].
+For video editing, we adopt the VBench Metrics.
 The evaluation criteria include Subject Consistency, Motion Smoothness, Aesthetic Quality, and Imaging Quality.
-Subject Consistency measures whether the subject (e.g., a person) remains consistent throughout the video by computing the similarity of DINO features [6] across frames.
-Motion Smoothness assesses the smoothness of motion in the generated video using motion priors from the video frame interpolation model [38].
-Aesthetic Quality evaluates the artistic and visual appeal of each frame as perceived by humans, leveraging the LAION aesthetic predictor [34].
-Imaging Quality examines the level of distortion in the generated frames (e.g., blurring or flickering) based on the MUSIQ image quality predictor [32].
-7 Stylization
-We provide more results on image stylization (Fig.
-9).
-In this task, we perform inversion on the image with a reference style, following the same overall pipeline as in image editing.
-RF-Edit demonstrates superior performance in stylized generation, effectively preserving the style of the source image while generating images that align with the user-provided prompts.
-8 Limitation
-Despite achieving impressive performance across various tasks, our methods occasionally exhibit performance instability, particularly in video editing.
-Additionally, the FLUX and OpenSora models contain a large number of parameters, leading to significant memory consumption, which may limit the applicability of our method on resource-constrained devices.
-To mitigate these issues, we plan to further optimize our method and explore its efficacy on more lightweight models.
+Subject Consistency measures whether the subject (e.g., a person) remains consistent throughout the video by computing the similarity of DINO features (Caron et al., 2021) across frames, which is similar to the CLIP Score for images.
+Motion Smoothness assesses the smoothness of motion in the generated video using motion priors from the video frame interpolation model (Li et al., 2023).
+Aesthetic Quality evaluates the artistic and visual appeal of each frame as perceived by humans, leveraging the LAION aesthetic predictor (LAION-AI, 2022).
+Imaging Quality examines the level of distortion in the generated frames (e.g., blurring or flickering) based on the MUSIQ image quality predictor (Ke et al., 2021).
+Appendix C Qualitative Results for Text-to-image Sampling
+Qualitative results for Text-to-image Sampling are shown in Figure 10.
+Experimental results demonstrate that compared to vanilla Rectified Flow, the RF-Solver sampler can generate higher-quality images that better align with human perception.
+Appendix D More Qualitative Results for Image Editing
+Here we provide more qualitative results for image editing and stylization (Figure 8).
+Appendix E More Qualitative Results for Video Editing
+HunyuanVideo (Kong et al., 2024) has recently demonstrated strong performance in text-to-video generation.
+The backbone of HunyuanVideo is similar to FLUX, which also contains several double-stream blocks, followed by single-stream blocks.
+We implement the RF-Solver and RF-Edit on HunyuanVideo, where the RF-Edit shares the feature in the single-stream block.
+The results are shown in Figure 9.
+Appendix F More Potential Applications
+RF-Solver is a universal sampler for rectified flow.
+Besides image and video editing, it is also potential on image or video generation (Xiao et al., 2025; Ma et al., 2024b, a; Lin et al., 2025) and other diffusion-based tasks (He et al., 2024; Fang et al., 2024; Guo et al., 2022).
+Furthermore, our proposed feature-sharing method in RF-Edit can also be applied in other image and video editing methods (Zhu et al., 2024b, a).